### v4

# MK-SGN: A Spiking Graph Convolutional Network with Multimodal Fusion and Knowledge Distillation for Skeleton-based Action Recognition

Abstract
In recent years, multimodal Graph Convolutional Networks (GCNs) have achieved remarkable performance in skeleton-based action recognition. The reliance on high-energy-consuming continuous floating-point operations inherent in GCN-based methods poses significant challenges for deployment in energy-constrained, battery-powered edge devices. To address these limitations, MK-SGN, a Spiking Graph Convolutional Network with Multimodal Fusion and Knowledge Distillation, is proposed to leverage the energy efficiency of Spiking Neural Networks (SNNs) for skeleton-based action recognition for the first time. By integrating the energy-saving properties of SNNs with the graph representation capabilities of GCNs, MK-SGN achieves significant reductions in energy consumption while maintaining competitive recognition accuracy. Firstly, we formulate a Spiking Multimodal Fusion (SMF) module to effectively fuse multimodal skeleton data represented as spike-form features. Secondly, we propose the Self-Attention Spiking Graph Convolution (SA-SGC) module and the Spiking Temporal Convolution (STC) module, to capture spatial relationships and temporal dynamics of spike-form features. Finally, we propose an integrated knowledge distillation strategy to transfer information from the multimodal GCN to the SGN, incorporating both intermediate-layer distillation and soft-label distillation to enhance the performance of the SGN. MK-SGN exhibits substantial advantages, surpassing state-of-the-art GCN frameworks in energy efficiency and outperforming state-of-the-art SNN frameworks in recognition accuracy. The proposed method achieves a remarkable reduction in energy consumption, exceeding 98% compared to conventional GCN-based approaches. This research establishes a robust baseline for developing high-performance, energy-efficient SNN-based models for skeleton-based action recognition
keywords:
Spiking Graph Convolution Network , Multimodal, Knowledge Distillation, Skeleton-based Action Recognition
[…]
[1] organization=Beijing Laboratory of Advanced Information Networks, Beijing Key Laboratory of Network System Architecture and Convergence, School of Information and Communication Engineering, Beijing University of Posts and Telecommunications, addressline=, city=Beijing, postcode=100876, state=, country=China
1 Introduction
Human action recognition is a critical task with wide-ranging applications in human-computer interaction, identity verification, and emergency detection [1]. Initial approaches predominantly relied on RGB-based methods that analyzed visual information extracted from video frames[2]. While effective in certain contexts, such methods encountered significant challenges in complex environments characterized by cluttered backgrounds, inadequate lighting, and occlusions, limiting their robustness and generalization capabilities.
Figure 1: Comparsion of (a) GCN and (b) SGN for skeleton-based Action Recognition. The former performs feature extraction and information propagation on nodes of the graph structure through graph convolutional layers, while the latter employs spike encoding to convert features into discrete spike-form, conducting feature extraction and information propagation in the time dimension of the spike-form feature.
To address limitations in earlier approaches, research has shifted towards skeleton-based action recognition, which utilizes 3D skeleton data obtained from depth sensors [3]. By isolating human action dynamics from environmental disturbances, skeleton-based recognition enhances robustness and reliability in action recognition tasks. Early methods employed Convolutional Neural Networks (CNNs) [4, 5, 6], Recurrent Neural Networks (RNNs) [8, 7], and Long Short-Term Memory networks (LSTMs) [9, 34] to model spatio-temporal features. However, limitations in representing the intrinsic graph structure of skeleton data prompted the development of Graph Convolutional Networks (GCNs), which represent a significant advancement in the field [10].
GCNs effectively model the spatial and temporal relationships within skeleton data, with the introduction of ST-GCN representing a key milestone [11]. The architecture employs spatial-temporal graph convolution to process skeletal data and laid the foundation for more advanced frameworks. 2S-AGCN improved performance by integrating a two-stream structure (Joint and Bone Streams) alongside adaptive graph convolution, enabling dynamic learning of skeletal topologies and refinement of temporal features [12]. Subsequent advancements aimed to balance efficiency and accuracy. Shift-GCN utilized a Channel-wise Shift Operation to simulate spatio-temporal information flow with minimal computational overhead [13], whereas CTR-GCN introduced channel-level topology optimization and attention mechanisms to enhance feature representation [14]. Building on foundational models, Info-GCN leverages information bottleneck theory for representation learning and extends the skeleton modality to six streams using the k-hop approach [15]. The development culminated in HULK, a multimodal large-scale model designed to unify diverse modalities within a robust framework. By harnessing its extensive capacity, HULK achieves state-of-the-art(SOTA) performance in precision and multimodal integration [16].
However, such performance in traditional methods often relies on deep network architectures typical of artificial neural networks (ANNs), which are heavily dependent on computationally expensive floating-point operations. Such reliance results in considerable computational energy demands. For example, lightweight models like Shift-GCN and CTR-GCN consume 46 mJ and 36.25 mJ per action sample, respectively, while 2S-AGCN requires up to 164.68 mJ per sample (single-sample energy consumption calculated based on Equation 37). Additionally, the pursuit of improved accuracy necessitates the integration of multimodal data, which requires separate training and testing for each modality, followed by ensemble fusion. These approaches significantly increase both computational complexity and energy consumption, posing substantial challenges for deployment on battery-powered edge devices and other energy-constrained environments
To address the high energy consumption and deployment challenges associated with ANN-based methods, Spiking Neural Networks (SNNs) provide a promising alternative, renowned for their energy efficiency [17]. Inspired by biomimetics, SNNs utilize event-driven mechanisms that enable exceptionally low energy consumption, making them particularly well-suited for resource-constrained environments. The energy-efficient design positions SNNs as a viable solution to the limitations of traditional ANN frameworks in practical applications [18, 19, 20]. SNNs have evolved significantly alongside advancements in ANNs, incorporating advanced architectures such as Spiking Graph Neural Networks [21, 22, 23], Spiking Recurrent Neural Networks [24, 25, 26], and Spiking Transformers [27, 28, 29]. Research efforts have primarily focused on tasks such as RGB image classification [51] and event-driven recognition [30, 31, 32], achieving notable success in reducing energy consumption while maintaining high accuracy.
Building on advancements in SNNs and GCNs, we propose a novel Spiking Graph Network (SGN) framework that, for the first time, integrates the energy efficiency of SNNs with the graph representation capabilities of GCNs for skeleton-based action recognition. Our work revolutionizes skeleton-based action recognition by establishing the first framework based on SNNs, unveiling their transformative potential in addressing diverse real-world challenges, particularly in environments with strict energy constraints.
The exploration of SGNs for skeleton action recognition is guided by several scientific challenges that prompt deep reflection:
  1. 1.
    Transformation of Skeleton Data into Spike-based Representations: What methodologies can effectively transform skeleton data, characterized by complex graphical structures, into spike-based representations compatible with the computational principles of SNNs?
  2. 2.
    Multimodal Feature Fusion within SNNs: What advanced techniques can facilitate the fusion of multimodal features in SNNs, ensuring consistency, coherence, and resilience in feature representation?
  3. 3.
    Compatibility with Vanilla GCNs: How can spike-based representations be effectively integrated with vanilla GCNs to ensure seamless functionality and robust performance in action recognition tasks?
  4. 4.
    Optimization of Energy Efficiency and Accuracy: What strategies can be employed to optimize the trade-off between energy consumption and recognition accuracy, achieving a balance that meets the requirements of energy-constrained environments?
The identified challenges emphasize crucial areas for innovation, driving progress at the intersection of SNNs and GCNs for achieving both energy-efficient and highly accurate action recognition. To address these issues, we first propose a Skeleton Spiking Coding (SSC) module to convert skeleton data into event spikes and a Spike-based Multimodal Fusion (SMF) module based on mutual information to compute the correlation between multimodal data for early fusion. The SMF module fully leverages the complementary information from multimodal skeleton data while avoiding the heavy training and testing burden associated with ensemble methods. Building on these components, we further develop a novel SGN block, which integrates Self-Attention-Spiking Graph Convolutions (SA-SGC) and Spiking Temporal Convolutions (STC). The SA-SGC employs a spiking attention mechanism to enhance the representation of spiking features, thereby improving classification performance. Meanwhile, the STC extracts temporal features from spike-form data derived from skeleton sequences, capturing fine-grained temporal dependencies to complement the spatial information encoded by the SA-SGC. This integrated design bridges the gap between SGN and traditional GCN in feature extraction and information propagation. Additionally, a GCN-to-SGN knowledge distillation method is formulated, comprising inner-layer feature distillation and soft label distillation. High-precision GCN models guide the training of SGN through these complementary strategies. A Feature Translation Module (FTM) is employed to convert inner-layer features from multimodal GCN into spike-form features, enabling effective inner-layer feature distillation. Soft label distillation transfers the output distribution knowledge from GCN to SGN, further enhancing model performance. The dual-distillation approach improves accuracy without increasing model complexity. Leveraging the maturity of GCN architectures and pre-trained weights, the process imposes no additional training burden. Ultimately, we propose Spiking Graph Convolutional Network with Multimodal Fusion and Knowledge Distillation (MK-SGN), the first framework to leverage SNNs for addressing skeleton-based action recognition. Extensive experiments validate the effectiveness of the proposed approach, demonstrating its ability to achieve energy-efficient and accurate performance while bridging the gap between SNNs and GCNs in this domain.
Overall, the main contributions of this paper can be summarized as follows:
(1) A novel MK-SGN framework is proposed, introducing skeleton-based action recognition into the domain of SNNs and establishing a foundation for future research in energy-efficient action recognition.
(2) An SMF module is formulated to fuse multimodal skeleton data in the form of spike-form features based on mutual information, effectively integrating complementary information while minimizing energy consumption and computational overhead. The novel SA-SGC and STC modules further enhance the representation of spatial relationships and temporal dynamics within the fused spike-form features.
(3) A knowledge distillation strategy is formulated to transfer knowledge from GCN to SGN, integrating both inner-layer feature distillation and soft label distillation for heterogeneous networks. To enable effective intermediate layer distillation, an FTM is designed to convert features from the intermediate layers of multimodal GCN into spike-form representations.
(4) Extensive experiments on three challenging datasets demonstrate the advantages of our work. The theoretical energy consumption of MK-SGN is 0.614 mJ per action sample. Compared to SOTA GCN-based methods, it reduces energy consumption by more than 98%. Moreover, we surpass the accuracy of SOTA SNN-based methods.
2 Related Work
2.1 Skeleton-based Action Recognition
In recent years, deep learning has significantly advanced skeleton-based action recognition [3]. Early research on skeleton data primarily uses CNNs to extract spatial features, focusing on patterns such as joint positions and spatial relationships [4, 5, 6]. RNNs, including the more advanced variant LSTMs, model temporal dynamics, capturing sequential dependencies in movement patterns over time. Such methods address spatial and temporal dimensions separately, with LSTMs excelling at learning long-term dependencies in motion sequences[8, 7, 34, 35, 36]. However, CNN- and RNN-based approaches fail to adequately exploit the graph-like structure of skeleton data, limiting their capacity to represent joint dependencies and hierarchical relationships fully.
Figure 2: The overview of the proposed network architecture. The teacher network consists of a 10-layer GCN that processes four multimodal skeleton datasets independently, followed by Global Average Pooling (GAP) and Fully Connected (FC) layers to generate soft labels. The student network is a 6-layer SGN that incorporates the following key modules: (a) The Spiking Coding Module (SSC), which transforms raw skeleton data into spike-form representations, enabling compatibility with spiking neural computations. (b) The Spiking Self-Attention Spiking Graph Convolution (SA-SGC) module, formulated to model the spatio-temporal dependencies within spike-form features effectively by combining self-attention and spiking graph convolution. (c) The Spiking Temporal Convolution (STC) module, which captures temporal patterns and refines temporal dynamics within spike-form features. Additionally, the Spike-based Multimodal Fusion Module (SMF) optimizes cross-modal integration using mutual information, and the Feature Transformation Module (FTM) maps inter-layer GCN features into spike-form representations for effective alignment with the SNN domain. Knowledge distillation is achieved through the FTM and supervision from the soft labels produced by the teacher network.
The introduction of ST-GCN marked a breakthrough by representing skeletons as graphs, enabling the explicit modeling of spatio-temporal relationships through graph convolutions[11]. Since then, numerous studies have aimed to improve GCN performance, focusing on both spatial and temporal modeling. In the spatial domain, methods have primarily focused on capturing intrinsic joint correlations through techniques such as learnable adjacency matrices [41, 42, 43], attention mechanisms [37, 38, 39, 40], and other adaptive strategies, to enhance spatial feature representations by dynamically modeling relationships between joints, thereby overcoming the limitations of static graph structures. In the temporal domain, methods have focused on improving the temporal receptive fields to effectively capture both long-term and short-term dependencies, employing multi-stage strategies to model complex action dynamics and achieve finer-grained temporal representations[14, 44]. The integration of multimodal skeleton data, including joint, bone, motion, and additional dynamic features, has significantly advanced the field. By leveraging complementary modalities, methods such as Shift-GCN, CTR-GCN, and Info-GCN enhance recognition performance [13, 14, 15]. The development culminated in Hulk [16], a multimodal large-scale model that unified diverse modalities into a comprehensive framework, achieving SOTA results. Despite advancements in skeleton-based action recognition, current methods rely on ANN-based architectures, which require significant floating-point operations and lead to high energy consumption. Although effective, scalability and energy efficiency remain challenges. SNNs present a potential solution but have yet to be explored in this context, leaving a notable gap in the research.
2.2 Spiking Neural Networks
Unlike ANNs, SNNs transmit information through discrete spike sequences, a mechanism inspired by biological principles. This spike-driven mode closely mimics brain function, offering significant potential for energy efficiency [17]. Early research in SNNs focused on various spiking neuron models designed to convert continuous inputs into spike sequences. Notable models include the Integrate-and-Fire (IF) neuron [45], which simplifies the process by accumulating input signals until a threshold is reached, and the Hodgkin-Huxley (HH) model [46], which provides a biophysically detailed representation of neuronal activity based on ion channel dynamics. The Leaky Integrate-and-Fire (LIF) neuron is widely regarded for its simplicity and efficiency [47]. The dynamics and update of the membrane potential are described as:
$\tau_{m}\frac{dV}{dt}=-\left(V-V_{\text{rest }}\right)+R\cdot I(t),$ (1)
$V(t+\Delta t)=\begin{cases}V_{\text{reset}},&\text{if }V(t)\geq V_{\text{thresh}}\\ V(t)+\left(\frac{-\left(V(t)-V_{\text{rest}}\right)+I(t)}{\tau_{m}}\right)\Delta t,&\text{otherwise}\end{cases},$ (2)
where $V(t)$ represents the membrane potential at the current time point, where $V_{\text{rest}}$ denotes the neuron’s resting potential. $I(t)$ indicates the input current. The membrane time constant is given by $\tau_{m}$, and $V_{\text{thresh}}$ is the threshold potential that triggers the firing of an action potential. After firing, the membrane potential is reset to $V_{\text{reset}}$. The model’s discrete update time step is $\Delta t$.
These advancements have led to the development of various spiking neural architectures, such as Spiking Graph Convolutional Networks [21, 22, 23], Spiking Transformers[27, 28, 29], Spiking Recurrent Neural Networks [24, 25, 26], and Spiking Convolutional Neural Networks [48, 49]. While existing research has primarily focused on traditional ANN tasks, significant progress has been made in spiking adaptations of Transformer models, particularly for RGB image processing. Studies in this area have demonstrated substantial energy savings without compromising accuracy[51]. Other research has focused on event-driven recognition, which is particularly well-suited for SNNs due to their ability to process sparse and asynchronous data [50]. Although SNNs have made significant progress in other areas, their application to skeleton-based action recognition remains underexplored. The event-driven nature of SNNs provides a promising approach to significantly improving energy efficiency, albeit with a certain trade-off in accuracy. Leveraging this potential could provide a viable solution to the challenge of balancing energy efficiency and accuracy in action recognition tasks.
3 Proposed Method
This section begins with the comprehensive architectural design of the MK-SGN, followed by an introduction to its Spiking Coding mechanism, the Spiking Multimodal Fusion Module, SGN Block and the Knowledge Distillation Method from GCN to SGN are then described in detail.
3.1 Overall Architecture
Our network architecture, illustrated in Figure 2, comprises a 10-layer multimodal teacher GCN network and a 6-layer SNN student network. The teacher GCN network processes the four multimodal skeleton datasets separately. The processed features from each modality are then passed through the Global Average Pooling (GAP) and Fully Connected (FC) layers to generate predictions. These predictions are finally fused to produce the soft labels.
The student SGN network processes each modality independently through the SSC module, transforming skeleton data into spike-form representations. Spike-form features are then fused via the SMF module, utilizing mutual information to optimize cross-modal integration. Each student network layer includes an SA-SGC module and a STC module, which are designed to capture and refine temporal and spatial features efficiently and effectively. Our approach implements knowledge distillation through inter-layer features and soft label distillation. The inter-layer features from the GCN teacher network are converted into spike-form representations via the FTM, ensuring alignment with the spiking domain and facilitating the training of the SNN student network. Soft labels produced by the teacher network provide supplementary supervisory signals, optimizing the learning process. The subsequent sections provide a detailed description of the design and functionality of the modules.
3.2 Spiking Coding Mechanism and Spiking Multimodal Fusion
This chapter discusses how to convert continuous real skeleton data into spiking-form representation. Then, we describe how to effectively fuse the multimodal spike-form features derived from multimodal skeleton data.
(a) Multimodal Skeleton Spiking Coding and Mutual Information Weight Matrix
(b) The detail structure of the SMIC Module
Figure 3: Multimodal Skeleton Spiking Coding and Mutual Information Weight Matrix Calculation Process
3.2.1 Skeleton Spiking Coding Module
For a given 3D skeleton sequence data $X_{i}\in\mathbb{R}^{C\times T\times V}$, the SSC Module first transforms it into an image-like representation. Then it expands it into a form suitable for spiking temporal representation illustrated in Figure 2(a). The process is described as follows:
$P_{0}=\operatorname{Expand}(X_{i})\in\mathbb{R}^{S\times C\times V\times T},$ (3)
$P_{i}=\operatorname{SN}(\operatorname{BN}(\operatorname{Conv2d}(P_{0})))\in\mathbb{R}^{S\times N\times V\times T}.$ (4)
In the initial step, the skeleton sequence data is transformed into an image-like representation to facilitate the structured feature extraction of spatial and temporal characteristics in subsequent stages. Subsequently, the $\operatorname{Expand}(\cdot)$ operation adds a spike time step dimension to the data, reorganizing the temporal structure to align with the requirements of spiking representation. The resulting expanded representation $P_{0}$ incorporates the additional spike time step dimension $S$ while preserving the original temporal dimension $T$, the graph structure $V$, and the channel size $C$.
[…]
3.2.2 Spike-based Multimodal Fusion Module
Multimodal skeleton data consists of four components: bones ($X_{\mathrm{B}}$), joints ($X_{\mathrm{J}}$), bone motions ($X_{\mathrm{BM}}$), and joint motions ($X_{\mathrm{JM}}$), each represented with dimensions $(C,T,V)$.
The SMF module is a mutual information-based fusion framework formulated specifically for the sparse spike-form features derived from multimodal data. In the SMF, the multimodal skeleton data are first encoded independently through the SSC module to obtain their respective spike-form features: $P_{\mathrm{B}},P_{\mathrm{J}},P_{\mathrm{BM}},P_{\mathrm{JM}}\in\mathbb{R}^{S\times D\times V\times T}$.
The spike-form features are then pairwise selected and concatenated along the feature dimension $D$ to construct a joint distribution. To construct a corresponding marginal distribution, one of the two spike-form features in each pair is randomly shuffled along the spike time step dimension $S$, and the same concatenation is performed. We propose the Spike-form Mutual Information Calculation (SMIC) network, formally constituted by an LSTM network, SN, FC, and GAP layers, to realize the spiking-driven mutual information prediction. The framework computes the mutual information between four sparse spike-form features and performs weighted fusion. By leveraging the computational efficiency of sparse spike-form features and maximizing the lower bound of mutual information, the SMIC network effectively captures shared information across modalities, facilitating both robust and energy-efficient multimodal feature fusion within SNN framework, as illustrated in Figure 3(b).
To illustrate the process, we take bone ($P_{\mathrm{B}}$) and joint ($P_{\mathrm{J}}$) data as an example. First, the spike-form features from the two modalities are concatenated along the feature dimension to construct a joint distribution:
$P_{\mathrm{Joint}}=\operatorname{cat}(P_{\mathrm{J}},P_{\mathrm{B}},\operatorname{dim}=1)\in\mathbb{R}^{S\times 2D\times V\times T}.$ (5)
To generate a marginal distribution, the spike-form feature $P_{\mathrm{B}}$ is shuffled randomly along the spike time step dimension ($S$), resulting in $P_{\mathrm{B}}^{\prime}$. The shuffled feature is then concatenated with $P_{\mathrm{J}}$:
$P_{\text{marginal}}=\operatorname{cat}(P_{\mathrm{J}},P_{\mathrm{B}}^{\prime},\operatorname{dim}=1)\in\mathbb{R}^{S\times 2D\times V\times T}.$ (6)
Both $P_{\mathrm{Joint}}$ and $P_{\text{marginal}}$ are then processed through an LSTM network equipped with SN layers. For the joint distribution, the output of the LSTM is passed through a FC layer followed by a GAP operation to compute:
$\boldsymbol{t}_{1}=\operatorname{SN}(\operatorname{LSTM}(P_{\mathrm{Joint}})),\quad\boldsymbol{t}=\operatorname{GAP}(\operatorname{FC}(\boldsymbol{t}_{1})).$ (7)
Similarly, for the marginal distribution, the output is processed through the same layers, and an additional exponential operation is applied:
$\boldsymbol{et}_{1}=\operatorname{SN}(\operatorname{LSTM}(P_{\text{marginal}})),\quad\boldsymbol{et}=\exp\left(\operatorname{GAP}(\operatorname{FC}(\boldsymbol{et}_{1}))\right).$ (8)
Using the predictions $\boldsymbol{t}$ and $\boldsymbol{et}$, the lower bound of mutual information is estimated as:
$mi_{-}lb=\bar{\boldsymbol{t}}-\log\overline{\boldsymbol{et}}.$ (9)
The calculation is performed pairwise for all modalities—bones, joints, bone motions, and joint motions—resulting in six pairwise mutual information values, which are organized into a symmetric matrix $\boldsymbol{MI}$ with zeros along the diagonal:
$\boldsymbol{MI}=\begin{bmatrix}0&\cdots&mi\_lb_{\mathrm{jb}}\\ \vdots&\ddots&\vdots\\ mi\_lb_{\mathrm{bj}}&\cdots&0\end{bmatrix}_{4\times 4}.$ (10)
To normalize the mutual information values, the sum along the rows of $\boldsymbol{MI}$ is computed, yielding the average mutual information weights:
$\boldsymbol{\overline{MI}}=\frac{\sum_{i=1}^{4}\boldsymbol{MI}_{i,:}}{\sum_{i=1}^{4}\sum_{j=1}^{4}\boldsymbol{MI}_{i,j}},$ (11)
$\boldsymbol{W}=\frac{\boldsymbol{\overline{MI}}-\min(\boldsymbol{\overline{MI}})}{\max(\boldsymbol{\overline{MI}})-\min(\boldsymbol{\overline{MI}})}\in\mathbb{R}^{4}.$ (12)
The fused spike-form features are then obtained by weighting the original features of all modalities using $\boldsymbol{W}$:
$P_{\text{fuse}}=\boldsymbol{W}*\left[P_{\mathrm{B}},P_{\mathrm{J}},P_{\mathrm{BM}},P_{\mathrm{JM}}\right]^{\boldsymbol{T}}\in\mathbb{R}^{S\times D\times V\times T}.$ (13)
Finally, the fused feature $P_{\text{fuse}}$ is fed into SGN Block for subsequent processing. By leveraging spiking-driven mutual information, the SMF module ensures the effective integration of sparse multimodal features, enhancing the robustness and accuracy of the overall action recognition framework.
3.3 SGN Block
Revisiting vanilla skeleton-based action recognition based on GCN, we construct our SGN using features converted into spike-form and adjacency matrices. The process can be summarized as follows:
$S=\widetilde{D}^{-\frac{1}{2}}\widetilde{A}\widetilde{D}^{-\frac{1}{2}},\quad$ (14)
where $S$ is the normalized adjacency matrix, $\widetilde{A}=A+I$ denotes the sum of the adjacency matrix $A$ and the identity matrix $I$ (representing self-connections). $\widetilde{D}$ is the diagonal degree matrix of $\widetilde{A}$. The graph convolutional output is then calculated using the following formula:
$H=S^{K}P\in\mathbb{R}^{S\times N\times T\times V},$ (15)
where, $H$ represents the output of the early spatial spiking graph convolution, $K$ is the number of graph convolutional layers, and $P$ is the spike-form feature input. The first layer of SGC can be expressed as:
$H^{(0)}=\mathrm{SN}(P)+\mathrm{SN}(HW_{sgc}^{(0)}),\quad$ (16)
where $W_{sgc}^{(0)}\in\mathbb{R}^{N^{(0)}\times N^{(1)}}$ is the learnable weight matrix of the first layer. Each subsequent layer has its learnable weights, denoted as $W_{sgc}^{(l)}\in\mathbb{R}^{N^{(l)}\times N^{(l+1)}}$.
To further enhance the model’s representation capabilities, we incorporate Self Spiking Attention (SSA) following the principles of existing spiking attention methods. These methods leverage a spiking-based $QKV$ structure with scaling factors $s$ to control large matrix multiplication values. The computation of SSA can be described as follows:
$Q^{(0)}=\mathrm{SN}_{Q}\left(\mathrm{BN}\left(H_{0}W^{(0)}_{Q}\right)\right),\quad$ (17)
$K^{(0)}=\mathrm{SN}_{K}\left(\mathrm{BN}\left(H_{0}W^{(0)}_{K}\right)\right),\quad$ (18)
$V^{(0)}=\mathrm{SN}_{V}\left(\mathrm{BN}\left(H_{0}W^{(0)}_{V}\right)\right),\quad$ (19)
$H^{(0)}_{SA}=H_{0}+\operatorname{SN}\left((Q^{(0)}{K^{(0)}}^{T})V^{(0)}\cdot s\right),\quad$ (20)
where, $Q^{(0)}$, $K^{(0)}$, and $V^{(0)}$ are computed through spiking neurons ($\mathrm{SN}$) after Batch Normalization ($\mathrm{BN}$) and linear transformations using learnable weights $W^{(0)}_{Q},W^{(0)}_{K},$ and $W^{(0)}_{V}$. The attention mechanism integrates the weighted interactions, while the scaling factor $s$ prevents large values in the spiking attention computation. $H^{(0)}_{SA}$ combines the residual connection $H_{0}$ with the attention-enhanced feature, defined as SA-SGC, illustrated in Figure 2(b), yielding a robust representation for spiking graph processing. After that, the first layer of STC can be written as:
$T^{(0)}=\operatorname{SN}\left(\mathrm{BN}\left(\operatorname{Conv2d}\left(H^{(0)}_{SA}\right)\right)\right)+\operatorname{SN}\left(H^{(0)}_{SA}\right).$ (21)
For subsequent layers, the spiking self attention-enhanced feature $H^{(l)}_{SA}$ is updated and combined with temporal features $T^{l-1}$, expressed as,illustrated in Figure 2(c):
$H^{(l)}_{SA}=\operatorname{SN}\left(T^{(l-1)}\right)+\operatorname{SN}\left(H^{(l-1)}_{SA}W_{sgc}^{(l-1)}\right),$ (22)
$T^{(l)}=\operatorname{SN}\left(\mathrm{BN}^{(l)}\left(\operatorname{Conv2d}^{(l)}\left(H^{(l)}_{SA}\right)\right)\right)+\operatorname{SN}\left(H^{(l)}_{SA}\right).$ (23)
Finally, the processed features are sent through GAP and a FC layer to output the prediction, as follows:
$y=\operatorname{FC}\left(\mathrm{GAP}\left(\operatorname{SN}\left(T^{(l)}\right)\right)\right),$ (24)
where $y\in\mathbb{R}^{U}$, and $U$ is the number of action classes. Based on this, a standard cross-entropy loss is adopted for model optimization, formulated as:
$L_{\mathrm{task}}=-\frac{1}{B}\sum_{b=1}^{B}\sum_{u=1}^{U}\mathrm{Y}_{bu}\log\mathrm{y}_{bu}+\left(1-\mathrm{Y}_{bu}\right)\log\left(1-\mathrm{y}_{bu}\right),$ (25)
where, $\mathrm{Y}$ is the ground-truth label, $\mathrm{y}$ is the predicted output, and $B$ denotes the number of training samples.
3.4 Knowledge Distillation of GCN-to-SGN
To leverage the powerful representation capabilities of multimodal GCNs, we propose a novel distillation framework with the multimodal GCN serving as the teacher network and the SGN as the student network, shown in Figure 4.
(a) Inner-layer feature distillation and soft label distillation.
(b) The detail structure of FTM
Figure 4: GCN-to-SGN Knowledge Distillation Method
The proposed approach fully exploits the rich knowledge embedded in pre-trained multimodal GCNs, transferring superior representational features to the SGN. By aligning the intermediate features and outputs of the SGN with those of the GCN, shown in Figure 4(a), the method significantly enhances SGN accuracy. The enhancement is achieved without introducing additional computational complexity or energy consumption, thereby preserving the inherent efficiency advantages of the SGN architecture. The distillation framework bridges the performance gap between GCNs and SGNs while maintaining the suitability of SGNs for energy-efficient applications.
3.4.1 Soft Label Distillation
In the proposed distillation framework, the CTR-GCN with a 4-stream architecture is utilized as the teacher network. Multimodal data inputs, including bone, joint, bone-motion, and joint-motion data, are processed independently by single-stream CTR-GCN. Single-modal soft labels, represented as $y_{b}$, $y_{j}$, $y_{bm}$, and $y_{jm}$, are generated corresponding to the respective modalities.
The outputs of the four modalities are then aggregated to form the multimodal soft labels $\hat{y}$, which serve as the knowledge to be distilled into the SGN. This aggregation is performed using weighted summation, as described below:
$\hat{y}=\alpha_{b}y_{b}+\alpha_{j}y_{j}+\alpha_{bm}y_{bm}+\alpha_{jm}y_{jm},$ (26)
where $\alpha_{b}$, $\alpha_{j}$, $\alpha_{bm}$, and $\alpha_{jm}$ are hyperparameters that control the relative contribution of each modality. These parameters are set based on the importance or reliability of the modalities during training.
The SGN then learns to approximate the multimodal teacher soft labels $\hat{y}$ through a regression-based loss function. The final learning objective for soft label distillation is expressed as:
$L_{sdk}=\|y-\hat{y}\|,$ (27)
where $y$ represents the predictions of the SGN. By aligning its outputs with the teacher’s multimodal soft labels, the SGN inherits the representational power of the multimodal GCN, improving its performance without increasing its energy consumption or computational cost.
3.4.2 Inner-layer feature distillation
The inner-layer features of heterogeneous GCN and SGN models differ significantly. While GCN typically processes continuous decimal values, SGN operates on sparse discrete spike sequences. To address this discrepancy, we propose the FTM, as illustrated in Figure 4(b). The FTM is an learnable module that transforms the continuous intermediate features of GCN into spike-form features, making them compatible with SGN.
The FTM concatenates multimodal features along the channel dimension and fuses them using depth-wise separable convolution. Subsequently, continuous features are expanded. Finally, spike-form features are generated through Conv2D, BN, and SN layers, ensuring compatibility between the teacher GCN and student SGN architectures while preserving the valuable structural and semantic information embedded in the GCN features.
In our distillation framework, the teacher GCN is composed of 10 GC-TC (Graph Convolution-Temporal Convolution) units, while the student SGN consists of 10 SGC-STC (Spiking Graph Convolution-Spiking Temporal Convolution) units. For feature distillation, we leverage intermediate-level features ($T_{\mathrm{Ab5}},T_{\mathrm{Aj5}},T_{\mathrm{Abm5}},T_{\mathrm{Ajm5}}$) and high-level features ($T_{\mathrm{Ab8}},T_{\mathrm{Aj8}},T_{\mathrm{Abm8}},T_{\mathrm{Ajm8}}$). The intermediate features are represented as: $T_{\mathrm{Ab5}},T_{\mathrm{Aj5}},T_{\mathrm{Abm5}},T_{\mathrm{Ajm5}}\in\mathbb{R}^{2C\times\frac{T}{2}\times V},$ while the high-level features are:$T_{\mathrm{Ab8}},T_{\mathrm{Aj8}},T_{\mathrm{Abm8}},T_{\mathrm{Ajm8}}\in\mathbb{R}^{4C\times\frac{T}{4}\times V}.$ These layers are selected because the expansion of channel dimensions and the changes in stride within the TCN network allow them to capture diverse temporal relationships and richer feature information. Taking the intermediate 5th layer as an example, the multimodal teacher features are first concatenated and fused as follows:
$T_{\mathrm{Am5}}=\operatorname{BN}\left(\operatorname{Conv2D}\left(\operatorname{cat}\left(T_{\mathrm{Ab5}},T_{\mathrm{Aj5}},T_{\mathrm{Abm5}},T_{\mathrm{Ajm5}}\right)\right)\right),$ (28)
where $T_{\mathrm{Am5}}\in\mathbb{R}^{8C\times\frac{T}{2}\times V}$ represents the fused multimodal teacher features. These features are then translated into spike-form features through the FTM:
$T_{\mathrm{Sm5}}=\operatorname{SN}\left(\operatorname{BN}\left(\operatorname{Conv2D}\left(\operatorname{Expand}\left(T_{\mathrm{Am5}}\right)\right)\right)\right),$ (29)
where $T_{\mathrm{Sm5}}\in\mathbb{R}^{S\times D\times V\times\frac{T}{2}}$ are the spike-form features. Similarly, the spike-form features for the high-level 8th layer ($T_{\mathrm{Sm8}}$) are calculated.
To align the student SGN features with the teacher GCN features, we use cosine similarity as the loss function. For the intermediate 5th-layer spike-form feature and the corresponding SGN 3rd-layer spike-form feature ($S_{3}$), the loss is defined as:
$L_{\mathrm{fkd}_{1}}=\frac{1}{J}\sum_{j=1}^{J}\left(1-\frac{T_{\mathrm{Sm5}}^{j}\cdot S_{3}^{j}}{\|T_{\mathrm{Sm5}}^{j}\|\cdot\|S_{3}^{j}\|}\right),$ (30)
where $J$ is the number of samples. Similarly, for the high-level 8th-layer spike-form feature and the SGN 5th-layer feature ($S_{5}$):
$L_{\mathrm{fkd}_{2}}=\frac{1}{J}\sum_{j=1}^{J}\left(1-\frac{T_{\mathrm{Sm8}}^{j}\cdot S_{5}^{j}}{\|T_{\mathrm{Sm8}}^{j}\|\cdot\|S_{5}^{j}\|}\right).$ (31)
The inner-layer feature distillation loss is then defined as:
$L_{\mathrm{fkd}}=\beta_{1}L_{\mathrm{fkd}_{1}}+\beta_{2}L_{\mathrm{fkd}_{2}},$ (32)
where $\beta_{1}$ and $\beta_{2}$ are hyperparameters that control the weighting of the intermediate and high-level distillation losses. This formulation ensures that the student network effectively learns from the teacher’s intermediate and high-level spike-form features while optimizing the FTM for feature translation.
Finally, the overall learning objective combines the inner-layer feature distillation loss with the task loss $L_{t}ask$ and the soft label distillation loss $L_{\mathrm{sdk}}$ as follows:
$L=\gamma_{1}L_{\mathrm{b}}+\gamma_{2}L_{\mathrm{sdk}}+\gamma_{3}L_{\mathrm{fkd}},$ (33)
where $\gamma_{1}$, $\gamma_{2}$, and $\gamma_{3}$ are hyperparameters that balance the contribution of each loss term. This unified objective ensures that the SGN effectively inherits the representational power of the GCN while maintaining its energy efficiency and computational advantages.
4 Experiments
In this chapter, we present the implementation details, ablation studies, result comparisons, and visualizations, respectively. The details will be described in the following subsections.
4.1 Experimental Setup
4.1.1 Datasets
We evaluate the proposed method on three widely used skeleton-based action recognition datasets: NTU-RGB+D, NTU-RGB+D 120, and NW-UCLA[52, 53, 54]. These datasets provide diverse action classes, extensive skeleton data, and varying levels of complexity, making them ideal benchmarks for comprehensive evaluation.
Table 1: The details of the datasets used in our experiments.
Dataset # Classes Split # Joints
Cross-Subject Cross-View (CSet)
# Train # Test # Train # Test
NTU RGB+D 60 40,320 16,560 37,920 18,960 25
NTU RGB+D 120 120 63,026 53,370 54,471 61,925 25
NW-UCLA 10 - - 1,039 455 20
NTU-RGB+D is a large-scale dataset comprising 60 action classes and 56,000 video clips. It includes detailed skeleton data covering a wide range of daily activities, captured using a Kinect v2 sensor. The dataset provides 3D coordinates of 25 major body joints, enabling precise modeling of human movements. NTU-RGB+D is widely recognized as a standard benchmark in the action recognition field due to its variety of actions and large-scale data volume. The dataset is evaluated under two standard settings: Cross-Subject (CSub), where training and testing samples are divided based on the subjects performing the actions, and Cross-View (CView), where training and testing samples are divided based on the camera views.
NTU-RGB+D 120, an extended version of NTU-RGB+D, significantly expands the dataset to include 120 action classes and over 114,000 video clips. This dataset offers richer skeleton data, making it one of the largest and most comprehensive resources for skeleton-based action recognition. Its inclusion of a broader range of actions and more extensive data volume provides an ideal foundation for developing and evaluating advanced action recognition algorithms. Similarly, this dataset is evaluated under two standard settings: Cross-Subject (CSub) and Cross-Setup (CSet), where training and testing samples are divided based on the camera setups used during data collection.
NW-UCLA is a smaller dataset consisting of 10 action classes and 1,494 video clips captured from multiple viewpoints. It provides 3D coordinates of 20 major body joints, offering diverse skeleton data for assessing the robustness of action recognition models to viewpoint variations. The dataset poses a challenge for models to perform consistently across different perspectives. The evaluation is conducted under the Cross-View setting, where training and testing samples are divided based on the camera angles used for data capture.
The specific statistics for the three datasets after the splits are summarized in Table 1.
4.1.2 Implementation Details
The experiments are conducted on a server equipped with an NVIDIA SMI V100 GPU with 30 GB of memory. The models are implemented using PyTorch[56] and SpikingJelly[55], which provide efficient frameworks for spiking neural networks. During training, we use two separate SGD optimizers: one dedicated to optimizing the SGN model and the other specifically designed to refine the FTM. This dual-optimizer strategy enhances the efficiency of the SGN and ensures effective feature transfer from the teacher to the student model.
The key hyperparameters for the LIF model used in all experiments are listed in Table 2. The training configurations for MK-SGN across different datasets are summarized in Table 3. Additionally, Table 4 details the hyperparameters for soft label aggregation and loss weighting, including the weights assigned to different modalities ($\alpha_{b},\alpha_{j},\alpha_{bm},\alpha_{jm}$) and the balancing factors for various loss components ($\beta_{1},\beta_{2},\gamma_{1},\gamma_{2},\gamma_{3}$).
Table 2: Hyperparameter settings for the LIF model.
Parameter Value
Threshold $V_{th}$ 1.0
Reset potential $V_{reset}$ 0
Decay factor $\tau$ 0.25
Surrogate function’s window size $a$ 1.0
The training configurations for MK-SGN across different datasets are summarized in Table 3. We adopt an initial learning rate of 0.1, a weight decay rate of 1e-4, and set the batch size based on the dataset: 64 for NTU-RGB+D 60, 32 for NTU-RGB+D 120, and 128 for NW-UCLA. Training is conducted for 64 epochs with a step decay schedule, reducing the learning rate after 50 epochs.
Table 3: Hyperparameter settings for MK-SGN training.
Parameter NTU-RGB+D NTU-RGB+D NW-UCLA
60 120
Learning Rate 0.1/0.1 0.1/0.1 0.1/0.1
Decay Rate 0.1/0.1 0.1/0.1 0.1/0.1
Batch Size 64 32 128
Time Steps 4 4 4
Training Epochs 64 64 64
Step 50 50 50
Dropout Rate 0.3 0.3 0.3
Weight Decay 1e-4 1e-4 1e-4
Optimizer SGD/SGD SGD/SGD SGD/SGD
Table 4: Hyperparameters for soft label aggregation and loss weighting.
Hyperparameter Description Value
$\alpha_{b}$ Weight for bone modality. 0.25
$\alpha_{j}$ Weight for joint modality. 0.25
$\alpha_{bm}$ Weight for bone motion modality. 0.25
$\alpha_{jm}$ Weight for joint motion modality. 0.25
$\beta_{1}$ Weight of intermediate distillation loss. 0.5
$\beta_{2}$ Weight of high-level distillation loss. 0.5
$\gamma_{1}$ Weight of task loss term. 1.0
$\gamma_{2}$ Weight of soft label loss term. 1.0
$\gamma_{3}$ Weight of feature distillation loss term. 1.0
In the testing phase, only the SGN model is evaluated, isolating it from the dual-optimizer framework used during training. This ensures that the SGN’s performance reflects its learned capabilities without additional computational overhead.
4.2 Ablation Studies
To verify the effectiveness and efficiency of the proposed model, we first replace the activation layers of the classic one-stream 2S-AGCN model with LIF neurons. This modified model, called Base-SGN*, is then trained and tested for comparison. Additionally, we compare our proposed model with Spikformer, a classic SNN-based transformer architecture. For a fair comparison, we adjust the layer structure and hidden dimension of Spikformer to match the MK-SGN architecture. These adjustments ensure that the comparison focuses on the modeling approach while minimizing the influence of structural differences. Based on these baselines, we conduct ablation and comparison experiments to evaluate the performance of MK-SGN.
4.2.1 Theoretical Energy Consumption Calculation
In the context of SNNs, synaptic operations (SOPs) play a critical role in determining theoretical energy consumption. As one of the key components of our experimental evaluation, we focus on calculating SOPs and the energy consumption of both Base-SGN* and MK-SGN, highlighting the efficiency of the proposed model.
Table 5: Evaluation on NTU-RGB+D. Param refers to the number of parameters. Power is the average theoretical energy consumption when predicting a skeleton from the NTU-RGB+D test set, whose calculation details are shown in Equation(35), Equation(36), and Equation(37).
Method Architecture Param Spike Time Step SMF Knowledge Distillation SOPs FLOPs Power Ntu-RGB+D 60
(M) (G) (G) ($mJ$) CSub(%) CView(%)
ANN (10-layer)ST-GCN 3.1 / / / / 3.48 16.01 81.5 88.3
SNN (10-layer)Base-SGN* 2.07 4 / / 0.60 7.17 0.536 64.2 71.3
(6-layer)SpikformerICLR 2023 4.78 4 / / 1.69 24.07 2.17 73.9 80.6
(6-layer) MK-SGN 1.89 4 / / 0.64 7.32 0.573 74.5 80.9
(6-layer) MK-SGN 1.89 4 / 0.66 7.32 0.582 75.9 82.1
(6-layer) MK-SGN 2.17 4 / 0.67 7.76 0.596 76.1 82.6
(6-layer) MK-SGN 2.17 4 0.68 7.76 0.614 78.5 85.6
  • 1.
    Note: FLOPs are calculated using the FVCORE library, which provides precise computation of floating-point operations.
Following [27, 28, 29], the firing rate $r$ of each LIF neuron, is defined as the firing probability per spike time step. Using the firing rate $r$, the synaptic operations for a given block or layer $l$ are expressed as:
$\operatorname{SOPs}(l)=r\times S\times\operatorname{FLOPs}(l),$ (34)
where $\operatorname{FLOPs}(l)$ refers to the floating-point operations of $l$, defined as the number of multiply-and-accumulate(MAC) operations performed in $l$ block or layer. And $\operatorname{SOPs}$ are the number of spike-based accumulate (AC) operations.
The theoretical energy consumption of Base-SGN and MK-SGN is calculated as follows:
$E_{\text{Base-SGN}}$ $=E_{\text{MAC}}\times\mathrm{FL}_{\mathrm{SNNConv}}^{1}$ (35)
$\quad+E_{\text{AC}}\times\left(\sum_{n=2}^{N}\mathrm{SOP}_{\mathrm{SNNConv}}^{n}+\sum_{m=1}^{M}\mathrm{SOP}_{\mathrm{SNNFC}}^{m}\right),$
$E_{\text{MK-SGN}}$ $=n_{m}\times E_{\text{MAC}}\times\mathrm{FL}_{\mathrm{SNNConv}}^{1}$ (36)
$\quad+E_{\text{AC}}\times\bigg{(}\frac{k\times(k-1)}{2}\mathrm{SOP}_{SMIC}+\sum_{n=n_{m}}^{N}\mathrm{SOP}_{\mathrm{SNNConv}}^{n}$
$\quad\quad+\sum_{m=1}^{M}\mathrm{SOP}_{\mathrm{SNNFC}}^{m}+\sum_{l=1}^{L}\mathrm{SOP}_{\mathrm{SSA}}^{l}\bigg{)}.$
In these equations, $\mathrm{FL}_{\mathrm{SNNConv}}^{1}$ represents the initial layer responsible for encoding skeleton data into spike-form, while $\mathrm{SOP}_{SMIC}$ accounts for the energy cost of modality fusion with $k$ denoting the number of modalities.
Energy consumption is estimated under the assumption of 45nm hardware, with $E_{\text{MAC}}=4.6\,\text{pJ}$ and $E_{\text{AC}}=0.9\,\text{pJ}$ [57]. The power consumption for a block $b$ in ANN (MAC-based) and SNN (SOP-based) frameworks is given by:
$\operatorname{Power}(b)=4.6\,\text{pJ}\times\operatorname{FLOPs}(b),$ (37)
$\operatorname{Power}(b)=0.9\,\text{pJ}\times\operatorname{SOPs}(b).$ (38)
These formulations demonstrate how MK-SGN balances high performance with reduced theoretical energy consumption. SOPs, representing synaptic operations, serve as a critical metric for evaluating the energy efficiency of spiking-driven architectures.
4.2.2 The Effect of Different Components
To analyze the energy efficiency of the proposed MK-SGN and evaluate the contribution of each component within the model, we use theoretical formulas to estimate the number of operations(OPs) required for classifying a single action sample. Specifically, we calculate the FLOPs for ANN-based models, as well as the FLOPs for SNN on traditional GPUs and the theoretical SOPs on neuromorphic hardware. Using these operation counts, we further compute the energy consumption per action sample during the testing phase.
Additionally, we evaluate the classification accuracy of MK-SGN through comprehensive ablation studies conducted on the NTU-RGB+D 60 dataset. The results, including energy efficiency and accuracy comparisons, are summarized in Table 5.
Base-SGN*
[…]
MK-SGN
Building on Base-SGN*, we compare its performance with the proposed MK-SGN. Even without knowledge distillation, the single-modality MK-SGN achieves significant accuracy improvements of 10.1% and 8.9% on XSub and XView settings, respectively, with only a minimal increase in OPs and theoretical power consumption. Compared to ST-GCN, although MK-SGN demonstrates a slight increase in FLOPs, this trade-off is offset by its extremely low SOPs value of 0.68, which highlights its suitability for deployment on neuromorphic hardware. Furthermore, MK-SGN reduces the parameter count by over 50% and decreases FLOPs and theoretical energy consumption by more than 70% compared to structurally similar models like Spikformer, while maintaining competitive computational efficiency.
Table 6: Performance comparison for different modality combinations using the SMF module. Accuracy is tested on Ntu-RGB+D Xsub
Architecture Modalities Parameters Ops Power Accuracy
(M) (G) (mJ) (%)
MK-SGN J 1.89 0.64 0.573 74.5
J+B 1.98 0.65 0.584 75.6
J+B+JM 2.08 0.66 0.591 75.9
J+B+JM+BM 2.17 0.67 0.596 76.1
Furthermore, MK-SGN achieves an accuracy improvement of 0.6% on XSub and 0.3% on XView compared to Spikformer, demonstrating its superior balance of performance and efficiency. These results highlight the effectiveness of MK-SGN in achieving high accuracy with significantly reduced computational and energy requirements.
SMF
In the case of late fusion in a multi-modal classification network, where $n_{n}$ represents the number of modalities, classifying a single action sample requires $n_{n}$ independent testing runs. This leads to a theoretical energy consumption scaled by $n_{n}$ times. Taking ST-GCN as an example, its theoretical OPs will become $4\times 3.48\,\text{G}$, and the Power will increase to $4\times 16.01\,\text{mJ}$.
To validate the effectiveness of the designed SMF module in avoiding the increased energy consumption caused by late fusion while simultaneously improving accuracy, we conduct more in-depth experiments, as shown in Table 6. Specifically, we incrementally integrate different modalities such as Joint (J), Joint+Bone (J+B), Joint+Bone+Joint-Motion (J+B+JM) and Joint+Bone+Joint-Motion +Bone-Motion(J+B+JM+BM)
[…]
These results validate that the SMF module effectively avoids the exponential increase in energy consumption caused by late fusion while consistently improving accuracy. Compared to single-modality fusion, the full multimodal fusion achieves an improvement of $1.6\%$ in accuracy for CSub, showcasing the robustness of the SMF module in efficiently integrating information across multiple modalities. This balance of efficiency and accuracy highlights the superiority of SMF over traditional late fusion methods.
Knowledge Distillation of GCN-to-SGN
Table 7: Accuracy comparison under different knowledge distillation settings. Accuracy is tested on Ntu-RGB+D Xsub
Setting SKD IFD Modalities Accuracy (%)
- - - - 74.5
Soft Label Only - J 75.1
Soft Label (Two Modal.) - J+B 75.4
Soft Label (Four Modal.) - J+B+JM+BM 75.6
Inner-layer (Low) - J+B+JM+BM 74.9
Inner-layer (High) - J+B+JM+BM 75.3
Full KD J+B+JM+BM 75.9
We compare the accuracy of SGN models trained with and without teacher guidance to evaluate the impact of our proposed knowledge distillation method. Knowledge distillation consists of two components: soft label distillation (SKD) and inner-layer feature distillation (IFD). To further analyze their contributions, we conduct experiments with and without each component. Additionally, we investigate the impact of soft label distillation using different modality combinations, as well as the effect of inner-layer feature distillation applied at different levels (low-level, high-level, or both).
The results in Table 7 demonstrate the significant contributions of SKD and IFD to the accuracy improvements of SGN models. With the introduction of SKD, accuracy improves steadily as more modalities are incorporated. Single-modality SKD achieves 75.1%, while combining two and four modalities further enhances the performance to 75.4% and 75.6%, respectively. This progression highlights the importance of multimodal knowledge transfer from the teacher network to the student model.
For IFD, focusing on low-level features yields an accuracy of 74.9%, while high-level features achieve 75.3%, indicating that the more abstract information encoded in high-level features plays a slightly greater role in guiding the student model. However, combining both low-level and high-level feature distillation demonstrates their complementary effects, achieving a final accuracy of 75.9% when combined with SKD in the full knowledge distillation (Full KD) setup.
The results highlight the importance of leveraging both output-level and feature-level guidance. While SKD aligns the student’s outputs with the teacher’s multimodal predictions, IFD bridges the gap in intermediate representations, allowing the student to better emulate the teacher’s internal structure. The findings validate the proposed knowledge distillation framework as an effective means of improving SGN model performance through both comprehensive guidance and multimodal integration.
Integrated Module Performance Comparison
We carefully evaluate the contributions of each component in the MK-SGN model, with results summarized in Table 5. The fully integrated MK-SGN achieves accuracies of 78.5% and 85.6% on XSub and XView settings, respectively, with a spike time step of 4. Compared to Spikformer, MK-SGN reduces the parameter count by 54.6% and FLOPs by 67.4%, significantly enhancing its feasibility for training on traditional GPUs. Moreover, MK-SGN decreases SOPs by 59.76% and theoretical energy consumption by 71.7%, while achieving accuracy improvements of 4.6% and 5.0%, demonstrating its ability to balance high performance with low energy consumption.
Compared to Base-SGN*, MK-SGN delivers substantial accuracy gains of 14.3% and 11.3% on XSub and XView, respectively. Additionally, when compared to the classic ANN-based model ST-GCN, MK-SGN achieves a remarkable 94.78% reduction in energy consumption while maintaining comparable accuracy. These results highlight the efficiency and effectiveness of MK-SGN, showcasing its suitability for energy-constrained applications without compromising performance.
4.3 Comparison with State-of-the-Art Methods
In this section, we conduct a comprehensive comparison between our proposed MK-SGN and SOTA methods for skeleton-based action recognition. The comparison is divided into two parts:(1). model architecture comparison and (2).metric comparison.
In the model architecture comparison, we focus on the design and layer configurations of MK-SGN in contrast to GCN and SNN models, using the NTU-RGB+D 60 dataset as a case study. This includes both publicly available results and those we reproduced under consistent experimental settings. The architecture analysis establishes the foundation for understanding the computational and operational differences, which are crucial for fair metric comparisons.
In the metric comparison, we evaluate MK-SGN against SOTA models across several dimensions, including accuracy, number of operations (OPs), and theoretical energy consumption. These metrics provide a quantitative perspective on MK-SGN’s ability to balance high performance with energy efficiency, particularly in energy-constrained environments.
By integrating both structural and performance analyses, we aim to demonstrate the superiority of MK-SGN in achieving competitive results with significantly reduced energy and computational overhead, establishing its practicality and effectiveness in real-world applications.
Table 8: Model’s architecture for NTU-RGB+D 60
Layer GCN GCN Structure Base-SGN MK-SGN SGN Structure Spikformer Spikformer Structure
(Input/Output) Layer Composition (Input/Output) (Input/Output) Layer Composition (Input/Output) Layer Composition
Fire layer - - 3/3 3/3 SC-(SMF) 3/256 SPS
1 3/64 GC-TC 3/64 3/64 SGC-STC 256/256 SSA-MLP
2 64/64 GC-TC 64/64 64/64 SGC-STC 256/256 SSA-MLP
3 64/64 GC-TC 64/64 64/128(stride=2) SGC-STC 256/256 SSA-MLP
4 64/64 GC-TC 64/64 128/128 SGC-STC 256/256 SSA-MLP
5 64/128(stride=2) GC-TC 64/128(stride=2) 128/256(stride=2) SGC-STC 256/256 SSA-MLP
6 128/128 GC-TC 128/128 256/256 SGC-STC 256/256 SSA-MLP
7 128/128 GC-TC 128/128 - SGC-STC - -
8 128/256(stride=2) GC-TC 128/256(stride=2) - SGC-STC - -
9 256/256 GC-TC 256/256 - SGC-STC - -
10 256/256 GC-TC 256/256 - SGC-STC - -
final 256/class number GAP-FC 256/class number 256/class number GAP-SN-FC 256/class number GAP-FC
4.3.1 Model Architecture Comparison
Table 8 summarizes the architectural differences between our proposed MK-SGN and various SOTA models, including GCN, and Spikformer, as applied to the NTU-RGB+D 60 dataset. This comparison highlights how each model is structured to handle skeleton-based action recognition tasks, with adaptations made to ensure fair performance evaluations.
The GCN model serves as a classical ANN-based approach, employing GC layers combined with TC layers for spatio-temporal feature extraction. Its structure consists of 10 layers of GC-TC operations, with progressively increasing feature dimensions and downsampling via strides, culminating in a GAP-FC layer for classification. This architecture represents a strong baseline for ANN-based recognition.
The Base-SGN* model acts as the baseline SNN structure, leveraging SGC and STC layers to process spike-form data. It follows a similar configuration to GCN but is adapted for spiking computation. However, it lacks the advanced features introduced in MK-SGN.
The Spikformer model, a spiking-based Transformer, incorporates SSA layers and MLP components for feature extraction. For this study, we modified Spikformer to align with MK-SGN in terms of layer count and hidden dimensions, ensuring a fair comparison of metrics such as energy efficiency and accuracy. This adjustment enables us to evaluate the performance of Spikformer in a setting directly comparable to MK-SGN. Similarly, other network variations like Spiking-driven-former, Spiking-driven-former v2 and Spiking Wavelet Transformer used in the study are adapted to maintain consistency across evaluations.
Finally, the MK-SGN model stands out with its innovative spiking-specific modules, including the SSC and SMF modules, which enhance multimodal integration and energy efficiency. The SGC-STC layers are optimized for spike-form data, enabling MK-SGN to balance high accuracy with reduced computational and energy requirements, setting it apart as an advanced solution for real-world applications.
This architectural comparison provides the groundwork for the subsequent metric analysis, ensuring that the reproduced models are aligned in structure and settings to allow for fair and meaningful performance evaluations.
4.3.2 Metric Comparison
Table 9: Comparative results on NTU RGB+D 60 and 120. We evaluate our model in terms of classification accuracy (%). X-Sub and X-View represent cross-subject and cross-view splits.
Method Architecture Backbone Spike Time Step Performance Accuracy (%)
Param OPs Power Csub Cview Csub CSet
(M) (G) (mJ) (60) (60) (120) (120)
Lie Group[58] ANN CNN / / / / 50.1 52.8 / /
HBRNN [8] ANN RNN / / / / 59.1 64.0 / /
Deep LSTM[52] ANN LSTM / / / / 60.7 67.3 / /
ST-GCN [11] ANN GCN / 3.1 3.48 16.01 81.5 88.3 70.7 73.2
2S-AGCN [12] ANN GCN / 3.48 35.8 164.68 88.5 95.0 82.5 84.2
Shift-GCN [13] ANN GCN / / 10 46 90.7 96.5 85.9 87.6
MS-G3D [59] ANN GCN / 3.19 67.63 311.09 91.5 96.2 86.9 88.4
CTR-GCN[14] ANN GCN / 1.46 7.88 36.25 92.4 96.8 88.9 90.6
Base-SGN* SNN SGN 4 2.07 0.60 0.536 64.2 71.3 45.2 47.3
Spikformer[27]ICLR 2023 SNN Spiking Transformer 4 4.78 1.69 2.17 73.9 80.1 61.7 63.7
Spike-driven Transformer[28] NIPS2023 SNN Spiking Transformer 4 4.77 1.57 1.93 73.4 80.6 62.3 64.1
Spike-driven Transformer V2[29]ICLR 2024 SNN Spiking Transformer 4 11.47 2.59 2.91 77.4 83.6 64.3 65.9
Spiking Wavelet Transformer[60]ECCV 2025 SNN Spiking Transformer 4 3.24 1.48 2.01 74.7 81.2 63.5 64.7
MK-SGN (Ours) SNN SGN 4 2.17 0.68 0.614 78.5 85.6 67.8 69.5
  • 1.
    Note: FLOPs are calculated using the FVCORE library, which provides precise computation of floating-point operations.
Table 10: Comparative results on NW-UCLA. We evaluate our model in terms of classification accuracy (%).
Method Architecture Backbone Times Param. OPs Power Acc
(M) (G) (mJ) (%)
Lie Group[58] ANN CNN / - - - 74.2
Actionlet ensemble [61] ANN CNN / - - - 76.0
HBRNN-L[63] ANN RNN / - - - 78.5
Ensemble TS-LSTM [62] ANN LSTM / - - - 89.2
Shift-GCN [13] ANN GCN / - 0.7 3.22 94.6
CTR-GCN [14] ANN GCN / 1.46 2.48 12.15 96.5
Base-SGN* SNN SGN 4 2.07 0.151 0.195 86.4
Spikformer[27]ICLR 2023 SNN Spiking Transformer 4 4.78 0.513 0.673 85.4
Spike-driven Transformer[28]NIPS2023 SNN Spiking Transformer 4 4.77 0.501 0.643 83.4
Spike-driven Transformer V2[29]ICLR 2024 SNN Spiking Transformer 4 11.47 0.754 0.914 89.4
Spiking Wavelet Transformer[60]ECCV 2025 SNN Spiking Transformer 4 3.24 0.451 0.697 86.9
MK-SGN (Ours) SNN SGN 4 2.17 0.165 0.207 92.3
  • 1.
    Note: FLOPs are calculated using the FVCORE library, which provides precise computation of floating-point operations.
We compare our proposed MK-SGN with the current SOTA ANNs and SNNs in skeleton-based action recognition, focusing on two critical performance metrics: OPs and theoretical energy consumption. The results are summarized in Table 9 and Table 10.
To ensure fairness, the reported results for existing multi-modal GCN methods utilize all four modalities: joint, bone, joint motion, and bone motion. For these methods, the softmax scores of multiple streams are fused by summing the scores from each stream. The OPs and theoretical energy consumption for these multi-modal methods are calculated by multiplying the values for a single stream by the total number of modalities.
Additionally, we reproduce the results of SOTA SNN models using only the joint modality to provide a direct and fair comparison with MK-SGN. These reproduced SNN models are carefully adapted to align with MK-SGN in terms of the number of layers and hidden dimensions. This ensures consistency in the experimental settings and allows for a robust evaluation of their efficiency and effectiveness.
Figure 5: Theoretical Energy Consumption Comparison
(a) Comparsion on NTU-RGB+D 60 Csub
(b) Comparsion on NTU-RGB+D 60 Cview
(c) Comparsion on NTU-RGB+D 120 Csub
(d) Comparsion on NTU-RGB+D 120 Cset
Figure 6: Comparsion of Accuracy with SOTA SNN
The comparison highlights that MK-SGN strikes a remarkable balance between accuracy and energy efficiency. Compared to existing methods, MK-SGN achieves significant reductions in theoretical energy consumption, emphasizing its suitability for energy-constrained scenarios, such as edge computing and real-time applications. By leveraging its spiking-driven architecture, MK-SGN demonstrates the ability to deliver competitive or superior performance while maintaining exceptional energy efficiency. Detailed results and analysis are provided in the following subsections.
Results on NTU-RGB+D 60 and 120
As shown in Table 9, MK-SGN demonstrates superior performance on both NTU-RGB+D 60 and NTU-RGB+D 120 datasets, showcasing its remarkable energy efficiency and competitive accuracy.
On the NTU-RGB+D 60 and NTU-RGB+D 120 datasets, MK-SGN achieves a significant reduction in theoretical energy consumption, requiring only 0.614 mJ per sample. This represents a 96.16% decrease compared to ST-GCN (16.01 mJ), a 99.63% decrease compared to 2S-AGCN (164.68 mJ), and a 99.80% decrease compared to MS-G3D (311.09 mJ). Even when compared to lightweight models like Shift-GCN (46 mJ) and CTR-GCN (36.25 mJ), MK-SGN reduces energy consumption by 98.67% and 98.31%, respectively, highlighting its remarkable efficiency advantage across all ANN-based GCN architectures, as showed in Figure 5.
Beyond energy savings, MK-SGN surpasses SOTA SNN-based models in both accuracy and efficiency, as highlighted in Table 9 and Figure 6. For NTU-RGB+D 60, MK-SGN achieves classification accuracies of 78.5% (Csub) and 85.6% (Cview), outperforming competitive SNN models such as Spikformer (73.9% Csub, 80.1% Cview) and Spike-driven Transformer V2 (77.4% Csub, 83.6% Cview). Importantly, MK-SGN achieves this with significantly reduced energy consumption, requiring only 0.614 mJ per sample compared to Spikformer’s 2.17 mJ and Spike-driven Transformer V2’s 2.91 mJ. MK-SGN also demonstrates an efficient parameter design with 2.17M parameters, far less than Spike-driven Transformer V2’s 11.47M, highlighting its lightweight and efficient architecture.
On the NTU-RGB+D 120 dataset, MK-SGN continues to outperform SNN models, achieving accuracies of 67.8% (Csub) and 69.5% (CSet). In comparison, Spikformer achieves 61.7% (Csub) and 63.7% (CSet), while Spike-driven Transformer V2 achieves 64.3% (Csub) and 65.9% (CSet). Despite these improvements in accuracy, MK-SGN remains highly energy-efficient, consuming only 0.614 mJ, compared to Spiking Wavelet Transformer’s 2.01 mJ and Spike-driven Transformer V2’s 2.91 mJ. This demonstrates MK-SGN’s superior performance and efficiency across datasets.
These results validate MK-SGN as a robust and energy-efficient solution for skeleton-based action recognition, excelling in scenarios requiring both high accuracy and low computational cost. By leveraging its spiking-driven architecture, MK-SGN sets a new benchmark for balancing performance and efficiency in SNN.
Results on NW-UCLA
On the NW-UCLA dataset, as summarized in Table 10, MK-SGN achieves a classification accuracy of 92.3%, surpassing SOTA SNN models. It outperforms models like Spikformer (85.4%) and Spike-driven Transformer V2 (89.4%) while maintaining significantly lower energy consumption, consuming only 0.207 mJ per sample compared to 0.914 mJ for Spike-driven Transformer V2.
This highlights MK-SGN’s ability to effectively balance accuracy and energy efficiency without compromising its computational performance. Its lightweight structure and optimized spiking modules allow for efficient recognition, particularly in scenarios where minimizing resource consumption is critical. These results further emphasize MK-SGN’s versatility and practicality across diverse application settings.
5 Conclusion
In this work, we propose the Spiking Graph Convolutional Network (MK-SGN), a novel architecture that combines the strengths of SNNs and graph convolutional networks GCNs for skeleton-based action recognition. This framework is, to the best of our knowledge, the first to effectively integrate spiking neural computation with graph-based modeling, offering a new perspective on energy-efficient and high-performance recognition methods.
The MK-SGN addresses the critical challenge of high energy consumption in traditional GCN-based methods while maintaining competitive performance. By leveraging multimodal fusion and knowledge distillation, MK-SGN not only optimizes the use of spiking neural computation but also enhances the ability to extract and integrate meaningful features across modalities. These innovations enable MK-SGN to significantly reduce energy consumption on three benchmark skeleton action recognition datasets, achieving over 98% reduction compared to the best ANN-based methods.
Moreover, MK-SGN consistently surpasses SOTA SNN models in recognition accuracy, achieving an average improvement of 5%. This highlights its ability to balance energy efficiency and computational effectiveness, making it particularly well-suited for real-time and resource-constrained applications. These results establish MK-SGN as a robust and practical solution for skeleton-based action recognition, paving the way for future exploration of spiking-based multimodal frameworks in other domains.

### v5

@@
 In recent years, multimodal Graph Convolutional Networks (GCNs) have achieved remarkable performance in skeleton-based action recognition.
-The reliance on high-energy-consuming continuous floating-point operations inherent in GCN-based methods poses significant challenges for deployment in energy-constrained, battery-powered edge devices.
-To address these limitations, MK-SGN, a Spiking Graph Convolutional Network with Multimodal Fusion and Knowledge Distillation, is proposed to leverage the energy efficiency of Spiking Neural Networks (SNNs) for skeleton-based action recognition for the first time.
-By integrating the energy-saving properties of SNNs with the graph representation capabilities of GCNs, MK-SGN achieves significant reductions in energy consumption while maintaining competitive recognition accuracy.
-Firstly, we formulate a Spiking Multimodal Fusion (SMF) module to effectively fuse multimodal skeleton data represented as spike-form features.
-Secondly, we propose the Self-Attention Spiking Graph Convolution (SA-SGC) module and the Spiking Temporal Convolution (STC) module, to capture spatial relationships and temporal dynamics of spike-form features.
-Finally, we propose an integrated knowledge distillation strategy to transfer information from the multimodal GCN to the SGN, incorporating both intermediate-layer distillation and soft-label distillation to enhance the performance of the SGN.
+The reliance on high-energy-consuming continuous floating-point operations inherent in GCN-based methods poses significant challenge for deployment in energy-constrained, battery-powered edge devices.
+To address the limitation, MK-SGN, a Spiking Graph Convolutional Network with Multimodal Fusion and Knowledge Distillation, is proposed to leverage the energy efficiency of Spiking Neural Networks (SNNs) for skeleton-based action recognition for the first time.
+By integrating the energy-saving properties of SNNs with the graph representation capabilities of GCNs, MK-SGN achieves significant reduction in energy consumption while maintaining competitive recognition accuracy.
+Firstly, we formulate a Spiking Multimodal Fusion (SMF) module to fuse the multimodal skeleton spike-form feature effectively.
+Secondly, we propose a Self-Attention Spiking Graph Convolution (SA-SGC) module and a Spiking Temporal Convolution (STC) module to capture spatial relationships and temporal dynamics of spike-form features, respectively.
+Finally, we propose an integrated knowledge distillation strategy to transfer information from the multimodal GCN to the SGN, incorporating inner-layer feature and soft-label distillation to enhance the SGN’s performance.
 MK-SGN exhibits substantial advantages, surpassing state-of-the-art GCN frameworks in energy efficiency and outperforming state-of-the-art SNN frameworks in recognition accuracy.
-The proposed method achieves a remarkable reduction in energy consumption, exceeding 98% compared to conventional GCN-based approaches.
-This research establishes a robust baseline for developing high-performance, energy-efficient SNN-based models for skeleton-based action recognition
+The proposed method remarkably reduces energy consumption, with 98% less consumed than conventional GCN-based approaches.
+In general, this research establishes a strong baseline for developing high-performance, energy-efficient SNN-based models for skeleton-based action recognition.
 keywords:
@@
 Initial approaches predominantly relied on RGB-based methods that analyzed visual information extracted from video frames[11].
-While effective in certain contexts, such methods encountered significant challenges in complex environments characterized by cluttered backgrounds, inadequate lighting, and occlusions, limiting their robustness and generalization capabilities.
-Figure 1: Comparsion of (a) GCN and (b) SGN for skeleton-based Action Recognition.
-The former performs feature extraction and information propagation on nodes of the graph structure through graph convolutional layers, while the latter employs spike encoding to convert features into discrete spike-form, conducting feature extraction and information propagation in the time dimension of the spike-form feature.
-To address limitations in earlier approaches, research has shifted towards skeleton-based action recognition, which utilizes 3D skeleton data obtained from depth sensors [3].
+While effective in certain contexts, such methods encountered significant challenges in complex environments characterized by cluttered backgrounds, inadequate illumination, and occlusions, limiting their robustness and generalization capabilities.
+Figure 1: Transformation from skeleton to spikes.
+The former GCN performs feature extraction and information propagation on nodes of the graph structure through graph convolutional layers, while the latter SGN employs spike encoding to convert features into discrete spike-form, conducting feature extraction and information propagation in the time dimension of the spike-form feature.
+Research has shifted towards skeleton-based action recognition to address limitations in these earlier approaches, which utilize 3D skeleton data obtained from depth sensors and pose estimation[12].
 By isolating human action dynamics from environmental disturbances, skeleton-based recognition enhances robustness and reliability in action recognition tasks.
 Early methods employed Convolutional Neural Networks (CNNs) [13, 14, 15], Recurrent Neural Networks (RNNs) [17, 16], and Long Short-Term Memory networks (LSTMs) [18, 43] to model spatio-temporal features.
-However, limitations in representing the intrinsic graph structure of skeleton data prompted the development of Graph Convolutional Networks (GCNs), which represent a significant advancement in the field [10].
+However, limitations in representing the intrinsic structure of skeleton data prompted the development of Graph Convolutional Networks (GCNs), which represent a significant advancement in the field [19].
 GCNs effectively model the spatial and temporal relationships within skeleton data, with the introduction of ST-GCN representing a key milestone [20].
-The architecture employs spatial-temporal graph convolution to process skeletal data and laid the foundation for more advanced frameworks.
-2S-AGCN improved performance by integrating a two-stream structure (Joint and Bone Streams) alongside adaptive graph convolution, enabling dynamic learning of skeletal topologies and refinement of temporal features [12].
-Subsequent advancements aimed to balance efficiency and accuracy.
-Shift-GCN utilized a Channel-wise Shift Operation to simulate spatio-temporal information flow with minimal computational overhead [13], whereas CTR-GCN introduced channel-level topology optimization and attention mechanisms to enhance feature representation [14].
-Building on foundational models, Info-GCN leverages information bottleneck theory for representation learning and extends the skeleton modality to six streams using the k-hop approach [15].
+The architecture employs spatial-temporal graph convolution to process skeleton data, laying the foundation for more advanced frameworks.
+2S-AGCN improves performance by integrating a two-stream structure (Joint and Bone Streams) alongside adaptive graph convolution, enabling dynamic learning of skeleton topologies and refinement of temporal features [21].
+Subsequent advancements aim to balance efficiency and accuracy.
+Shift-GCN utilizes a channel-wise shift operation to simulate spatio-temporal information flow with lightweight computational overhead[22].
+CTR-GCN introduces channel-level topology optimization and attention mechanisms to enhance feature representation [23].
+Building on these, Info-GCN leverages information bottleneck theory for representation learning and extends the skeleton modality to six streams using the k-hop approach [24].
 The development culminated in HULK, a multimodal large-scale model designed to unify diverse modalities within a robust framework.
 By harnessing its extensive capacity, HULK achieves state-of-the-art(SOTA) performance in precision and multimodal integration [25].
-However, such performance in traditional methods often relies on deep network architectures typical of artificial neural networks (ANNs), which are heavily dependent on computationally expensive floating-point operations.
+However, such methods rely on deep network architectures typical of artificial neural networks (ANNs), which depend heavily on computationally expensive floating-point operations.
 Such reliance results in considerable computational energy demands.
-For example, lightweight models like Shift-GCN and CTR-GCN consume 46 mJ and 36.25 mJ per action sample, respectively, while 2S-AGCN requires up to 164.68 mJ per sample (single-sample energy consumption calculated based on Equation 37).
-Additionally, the pursuit of improved accuracy necessitates the integration of multimodal data, which requires separate training and testing for each modality, followed by ensemble fusion.
-These approaches significantly increase both computational complexity and energy consumption, posing substantial challenges for deployment on battery-powered edge devices and other energy-constrained environments
+For example, lightweight models like Shift-GCN and CTR-GCN consume 46 mJ and 36.25 mJ per action sample, respectively, while 2S-AGCN requires up to 182.87 mJ per sample (single-sample energy consumption calculated based on Equation 36).
+Additionally, the pursuit of better performance necessitates the integration of multimodal data, which requires separate training and testing for each modality, followed by ensemble fusion.
+These characteristics of ANN-based methods lead to significant computational complexity and energy consumption, posing substantial challenges for deployment on battery-powered edge devices and other energy-constrained environments.
 To address the high energy consumption and deployment challenges associated with ANN-based methods, Spiking Neural Networks (SNNs) provide a promising alternative, renowned for their energy efficiency [26].
@@
 Building on advancements in SNNs and GCNs, we propose a novel Spiking Graph Network (SGN) framework that, for the first time, integrates the energy efficiency of SNNs with the graph representation capabilities of GCNs for skeleton-based action recognition.
-Our work revolutionizes skeleton-based action recognition by establishing the first framework based on SNNs, unveiling their transformative potential in addressing diverse real-world challenges, particularly in environments with strict energy constraints.
+Our work establishes the first framework leveraging SNNs for skeleton-based action recognition, significantly advancing this task’s potential in energy-constrained environments.
 The exploration of SGNs for skeleton action recognition is guided by several scientific challenges that prompt deep reflection:
   1. 1.
-    Transformation of Skeleton Data into Spike-based Representations: What methodologies can effectively transform skeleton data, characterized by complex graphical structures, into spike-based representations compatible with the computational principles of SNNs?
+    Transformation of Skeleton Data into Spike-based Representations: Effectively transforming skeleton data, characterized by complex graphical structures, into spike-based representations compatible with the computational principles of SNNs remains a crucial challenge.
   2. 2.
-    Multimodal Feature Fusion within SNNs: What advanced techniques can facilitate the fusion of multimodal features in SNNs, ensuring consistency, coherence, and resilience in feature representation?
+    Multimodal Feature Fusion within SNNs: Facilitating the fusion of multimodal features in SNNs requires advanced methods to ensure consistency, coherence, and resilience in feature representation.
   3. 3.
-    Compatibility with Vanilla GCNs: How can spike-based representations be effectively integrated with vanilla GCNs to ensure seamless functionality and robust performance in action recognition tasks?
+    Compatibility with Vanilla GCNs: Adapting the vanilla GCN framework to spike-based representations while ensuring robust performance in action recognition poses a significant research challenge.
   4. 4.
-    Optimization of Energy Efficiency and Accuracy: What strategies can be employed to optimize the trade-off between energy consumption and recognition accuracy, achieving a balance that meets the requirements of energy-constrained environments?
-The identified challenges emphasize crucial areas for innovation, driving progress at the intersection of SNNs and GCNs for achieving both energy-efficient and highly accurate action recognition.
+    Optimization of Energy Efficiency and Accuracy: Balancing energy consumption and recognition accuracy to meet the requirements of energy-constrained environments remains an essential goal.
 To address these issues, we first propose a Skeleton Spiking Coding (SSC) module to convert skeleton data into event spikes and a Spike-based Multimodal Fusion (SMF) module based on mutual information to compute the correlation between multimodal data for early fusion.
 The SMF module fully leverages the complementary information from multimodal skeleton data while avoiding the heavy training and testing burden associated with ensemble methods.
-Building on these components, we further develop a novel SGN block, which integrates Self-Attention-Spiking Graph Convolutions (SA-SGC) and Spiking Temporal Convolutions (STC).
-The SA-SGC employs a spiking attention mechanism to enhance the representation of spiking features, thereby improving classification performance.
-Meanwhile, the STC extracts temporal features from spike-form data derived from skeleton sequences, capturing fine-grained temporal dependencies to complement the spatial information encoded by the SA-SGC.
-This integrated design bridges the gap between SGN and traditional GCN in feature extraction and information propagation.
+Based on these components, we propose a novel SGN block integrating a Self-Attention-Spiking Graph Convolutions (SA-SGC) module and a Spiking Temporal Convolutions (STC) module.
+The SA-SGC captures the spatial dependencies of skeleton sequences represented in spiking form through a graph topology and enhances them via a spiking attention mechanism, while the STC extracts temporal dependencies to complement the spatial information represented by the SA-SGC.
+This integrated design provides a unified framework that leverages the strengths of SGN and vanilla GCNs for effective feature extraction and information propagation.
 Additionally, a GCN-to-SGN knowledge distillation method is formulated, comprising inner-layer feature distillation and soft label distillation.
 High-precision GCN models guide the training of SGN through these complementary strategies.
-A Feature Translation Module (FTM) is employed to convert inner-layer features from multimodal GCN into spike-form features, enabling effective inner-layer feature distillation.
+A Feature Translation Module (FTM) is formulated to convert inner-layer features from multimodal GCN into spike-form features, enabling effective inner-layer feature distillation.
 Soft label distillation transfers the output distribution knowledge from GCN to SGN, further enhancing model performance.
@@
 Leveraging the maturity of GCN architectures and pre-trained weights, the process imposes no additional training burden.
-Ultimately, we propose Spiking Graph Convolutional Network with Multimodal Fusion and Knowledge Distillation (MK-SGN), the first framework to leverage SNNs for addressing skeleton-based action recognition.
+Ultimately, we propose Spiking Graph Convolutional Network with Multimodal Fusion and Knowledge Distillation (MK-SGN), the first framework to leverage SNNs for skeleton-based action recognition.
 Extensive experiments validate the effectiveness of the proposed approach, demonstrating its ability to achieve energy-efficient and accurate performance while bridging the gap between SNNs and GCNs in this domain.
@@
 (1) A novel MK-SGN framework is proposed, introducing skeleton-based action recognition into the domain of SNNs and establishing a foundation for future research in energy-efficient action recognition.
-(2) An SMF module is formulated to fuse multimodal skeleton data in the form of spike-form features based on mutual information, effectively integrating complementary information while minimizing energy consumption and computational overhead.
-The novel SA-SGC and STC modules further enhance the representation of spatial relationships and temporal dynamics within the fused spike-form features.
-(3) A knowledge distillation strategy is formulated to transfer knowledge from GCN to SGN, integrating both inner-layer feature distillation and soft label distillation for heterogeneous networks.
-To enable effective intermediate layer distillation, an FTM is designed to convert features from the intermediate layers of multimodal GCN into spike-form representations.
+(2)An SMF module is formulated to fuse multimodal skeleton data in spike form based on mutual information, effectively integrating complementary information while reducing energy consumption and computational overhead.
+The novel SA-SGC and STC modules further enhance the representation of spatial relationships and temporal dynamics within the fused spike-form features, respectively.
+(3)A knowledge distillation strategy is established to transfer knowledge from GCN to SGN, integrating inner-layer feature and soft-label distillation for heterogeneous networks.
+To enable effective inner-layer feature distillation, an FTM is employed to convert features from the intermediate layers of multimodal GCN into spike-form representations.
 (4) Extensive experiments on three challenging datasets demonstrate the advantages of our work.
-The theoretical energy consumption of MK-SGN is 0.614 mJ per action sample.
-Compared to SOTA GCN-based methods, it reduces energy consumption by more than 98%.
+The theoretical energy consumption of MK-SGN is 0.596 mJ per action sample.
+In general, Compared to SOTA GCN-based methods, with 98% less energy consumed than conventional GCN-based approaches.
 Moreover, we surpass the accuracy of SOTA SNN-based methods.
@@
 In recent years, deep learning has significantly advanced skeleton-based action recognition [12].
-Early research on skeleton data primarily uses CNNs to extract spatial features, focusing on patterns such as joint positions and spatial relationships [4, 5, 6].
-RNNs, including the more advanced variant LSTMs, model temporal dynamics, capturing sequential dependencies in movement patterns over time.
-Such methods address spatial and temporal dimensions separately, with LSTMs excelling at learning long-term dependencies in motion sequences[8, 7, 34, 35, 36].
-However, CNN- and RNN-based approaches fail to adequately exploit the graph-like structure of skeleton data, limiting their capacity to represent joint dependencies and hierarchical relationships fully.
-Figure 2: The overview of the proposed network architecture.
-The teacher network consists of a 10-layer GCN that processes four multimodal skeleton datasets independently, followed by Global Average Pooling (GAP) and Fully Connected (FC) layers to generate soft labels.
-The student network is a 6-layer SGN that incorporates the following key modules: (a) The Spiking Coding Module (SSC), which transforms raw skeleton data into spike-form representations, enabling compatibility with spiking neural computations.
-(b) The Spiking Self-Attention Spiking Graph Convolution (SA-SGC) module, formulated to model the spatio-temporal dependencies within spike-form features effectively by combining self-attention and spiking graph convolution.
-(c) The Spiking Temporal Convolution (STC) module, which captures temporal patterns and refines temporal dynamics within spike-form features.
-Additionally, the Spike-based Multimodal Fusion Module (SMF) optimizes cross-modal integration using mutual information, and the Feature Transformation Module (FTM) maps inter-layer GCN features into spike-form representations for effective alignment with the SNN domain.
-Knowledge distillation is achieved through the FTM and supervision from the soft labels produced by the teacher network.
-The introduction of ST-GCN marked a breakthrough by representing skeletons as graphs, enabling the explicit modeling of spatio-temporal relationships through graph convolutions[11].
-Since then, numerous studies have aimed to improve GCN performance, focusing on both spatial and temporal modeling.
-In the spatial domain, methods have primarily focused on capturing intrinsic joint correlations through techniques such as learnable adjacency matrices [41, 42, 43], attention mechanisms [37, 38, 39, 40], and other adaptive strategies, to enhance spatial feature representations by dynamically modeling relationships between joints, thereby overcoming the limitations of static graph structures.
-In the temporal domain, methods have focused on improving the temporal receptive fields to effectively capture both long-term and short-term dependencies, employing multi-stage strategies to model complex action dynamics and achieve finer-grained temporal representations[14, 44].
-The integration of multimodal skeleton data, including joint, bone, motion, and additional dynamic features, has significantly advanced the field.
-By leveraging complementary modalities, methods such as Shift-GCN, CTR-GCN, and Info-GCN enhance recognition performance [13, 14, 15].
-The development culminated in Hulk [16], a multimodal large-scale model that unified diverse modalities into a comprehensive framework, achieving SOTA results.
-Despite advancements in skeleton-based action recognition, current methods rely on ANN-based architectures, which require significant floating-point operations and lead to high energy consumption.
-Although effective, scalability and energy efficiency remain challenges.
-SNNs present a potential solution but have yet to be explored in this context, leaving a notable gap in the research.
+Early research primarily employed CNNs to extract spatial features, capturing joint positions and local spatial relationships [13, 14, 15].
+RNNs, especially LSTMs, were then applied to model temporal dynamics, excelling at capturing long-term sequential dependencies in motion data [17, 16, 43, 44, 45].
+While these approaches successfully addressed spatial and temporal dimensions separately, they failed to fully exploit the inherent graph structure of skeleton data, limiting their capacity to represent complex joint dependencies and hierarchical relationships.
+The introduction of ST-GCN [20] marked a turning point by representing skeletons as graphs and applying graph convolutions to jointly capture spatial and temporal patterns.
+However, its reliance on a fixed adjacency matrix restricted adaptability to dynamic joint interactions.
+To overcome this limitation, subsequent studies proposed learnable adjacency matrices [50, 51, 52], which allow edges to be optimized for specific actions, though at the cost of higher model complexity and potential overfitting.
+Another line of work incorporates attention mechanisms [46, 47, 48, 49], which selectively emphasize informative joints and capture long-range dependencies, thereby enriching feature representations.
+Nevertheless, these methods often increase computational cost.
+Other adaptive strategies, such as multi-scale sampling or channel-mixing, further improve robustness and flexibility, but the balance among accuracy, efficiency, and generalizability remains an open challenge.
+Beyond spatial modeling, researchers have also enhanced temporal modeling.
+MS-G3D [68] introduced multi-scale temporal modeling through dilated temporal windows and cross-spacetime graph convolutions, improving the capture of long-range temporal dependencies, while STA-GCN [53] used temporal adaptive graph convolutions to jointly model short- and long-term patterns.Together, these advances demonstrate the importance of refining both spatial and temporal receptive fields to improve overall recognition performance.
+More recently, multimodal integration has become a key driver of progress.
+Approaches that incorporate joint, bone, motion, and other dynamic cues enrich feature diversity.
+With the development of multimodal foundation models, Hulk [25] unified heterogeneous skeleton modalities into a comprehensive framework, achieving state-of-the-art accuracy.
+BlockGCN [7] dynamically redefined topology awareness to better capture dependencies, offering accuracy comparable to Hulk while retaining lower computational cost.
+This contrast highlights the trade-off between large-scale multimodal integration and efficient graph-based modeling.
+The rise of large language models (LLMs) has also inspired novel directions.
+PURLS [8] leveraged GPT-3 to enhance zero-shot action recognition by generating semantic action descriptions, facilitating transfer to unseen classes.
+Neuron [1] advanced this line with evolving micro-prototypes for tighter alignment between skeleton data and semantic context.
+Similarly, CrossGLG [2] applied a global–local–global LLM-guided strategy to one-shot recognition, enriching feature abstraction across multiple levels.
+Although not directly targeting action recognition, M-R²ET [6] explored motion retargeting through decoupled skeleton and shape modeling, offering valuable insights into handling diverse skeleton structures and transfer learning.
+Despite these advances, most existing methods remain grounded in conventional ANNs, which rely heavily on floating-point operations and thus incur high computational and energy costs.
+This limits their scalability and deployment in resource-constrained scenarios.
+SNNs, known for their event-driven and energy-efficient computation, present a promising alternative.
+However, their potential in skeleton-based action recognition has not yet been explored, leaving a critical research gap that our work aims to address.
 2.2 Spiking Neural Networks
@@
 Early research in SNNs focused on various spiking neuron models designed to convert continuous inputs into spike sequences.
-Notable models include the Integrate-and-Fire (IF) neuron [45], which simplifies the process by accumulating input signals until a threshold is reached, and the Hodgkin-Huxley (HH) model [46], which provides a biophysically detailed representation of neuronal activity based on ion channel dynamics.
+Notable models include the Integrate-and-Fire (IF) neuron [54], which simplifies the process by accumulating input signals until a threshold is reached.
+The Hodgkin-Huxley (HH) model [55] provides a biophysically detailed representation of neuronal activity based on ion channel dynamics.
 The Leaky Integrate-and-Fire (LIF) neuron is widely regarded for its simplicity and efficiency [56].
-The dynamics and update of the membrane potential are described as:
+The dynamics and update of the membrane potential are described as follows:
 $\tau_{m}\frac{dV}{dt}=-\left(V-V_{\text{rest }}\right)+R\cdot I(t),$ (1)
-$V(t+\Delta t)=\begin{cases}V_{\text{reset}},&\text{if }V(t)\geq V_{\text{thresh}}\\ V(t)+\left(\frac{-\left(V(t)-V_{\text{rest}}\right)+I(t)}{\tau_{m}}\right)\Delta t,&\text{otherwise}\end{cases},$ (2)
+$V(t+\Delta t)=\begin{cases}V_{\text{reset}},&\text{\hskip-20.00003ptif }V(t)\geq V_{\text{thresh}}\\ V(t)+\left(\frac{-\left(V(t)-V_{\text{rest}}\right)+I(t)}{\tau_{m}}\right)\Delta t,&\text{otherwise}\end{cases},$ (2)
 where $V(t)$ represents the membrane potential at the current time point, where $V_{\text{rest}}$ denotes the neuron’s resting potential.
@@
 The model’s discrete update time step is $\Delta t$.
-These advancements have led to the development of various spiking neural architectures, such as Spiking Graph Convolutional Networks [21, 22, 23], Spiking Transformers[27, 28, 29], Spiking Recurrent Neural Networks [24, 25, 26], and Spiking Convolutional Neural Networks [48, 49].
-While existing research has primarily focused on traditional ANN tasks, significant progress has been made in spiking adaptations of Transformer models, particularly for RGB image processing.
+These advancements have led to the development of various spiking neural architectures, such as Spiking Graph Convolutional Networks [30, 31, 32], which integrate Spiking Neural Networks into Graph Convolutional Networks to efficiently process dynamic graph data while reducing computational costs.
+In the domain of sequence processing, Spiking Transformers [36, 37, 38] leverage spiking self-attention mechanisms to replace traditional matrix multiplication with sparse operations, improving energy efficiency in tasks like RGB image processing.
+Meanwhile, Spiking Recurrent Neural Networks[33, 34, 35] combine the temporal dynamics of recurrent networks with spiking neuron models, offering superior performance in sequential tasks such as speech and gesture recognition, while significantly reducing power consumption.
+Lastly, Spiking Convolutional Neural Networks [57, 58] adapt spiking neurons to convolutional layers, enabling efficient processing of spatial correlations in image-based tasks, providing an energy-efficient alternative for real-time applications on neuromorphic hardware.
 Studies in this area have demonstrated substantial energy savings without compromising accuracy[60].
-Other research has focused on event-driven recognition, which is particularly well-suited for SNNs due to their ability to process sparse and asynchronous data [50].
+Other research has focused on event-driven recognition, leveraging SNNs for their suitability in processing sparse and asynchronous data [59].
 Although SNNs have made significant progress in other areas, their application to skeleton-based action recognition remains underexplored.
-The event-driven nature of SNNs provides a promising approach to significantly improving energy efficiency, albeit with a certain trade-off in accuracy.
-Leveraging this potential could provide a viable solution to the challenge of balancing energy efficiency and accuracy in action recognition tasks.
+The event-driven nature of SNNs provides a promising approach to significantly improving energy efficiency, albeit with an inevitable trade-off in accuracy.
+Leveraging this potential could provide a viable solution to balancing energy efficiency and accuracy in action recognition tasks.
 3 Proposed Method
-This section begins with the comprehensive architectural design of the MK-SGN, followed by an introduction to its Spiking Coding mechanism, the Spiking Multimodal Fusion Module, SGN Block and the Knowledge Distillation Method from GCN to SGN are then described in detail.
+This section provides an overview of MK-SGN’s comprehensive architecture, followed by a detailed explanation of its Skeleton Spiking Coding module and Spiking Multimodal Fusion Module.
+Additionally, the SGN Block and the Knowledge Distillation Method from GCN to SGN are thoroughly discussed.
+Figure 2: The overview of the proposed MK-SGN architecture.
+The teacher network comprises a 10-layer GCN that processes four multimodal skeleton data separately, followed by Global Average Pooling (GAP) and Fully Connected (FC) layers to generate soft labels.
+The student network is a 6-layer SGN that incorporates the following key modules: (a) The Skeleton Spiking Coding (SSC) module, responsible for transforming raw skeleton data into spike-form feature to ensure compatibility with spiking neural computations.
+(b) The Spiking Self-Attention Spiking Graph Convolution (SA-SGC) module is formulated to model spatial dependencies and enhance feature extraction by integrating Spiking Graph Convolution and Spiking Self-Attention (SSA).
+(c) The Spiking Temporal Convolution (STC) module is proposed to capture temporal patterns and refine temporal dynamics within spike-form features.
+Additionally, the Spike-based Multimodal Fusion Module (SMF) optimizes cross-modal integration based on mutual information, and the Feature Transformation Module (FTM) maps inter-layer GCN features into spike-form features for effective alignment with the SGN.
+Knowledge distillation is achieved through the FTM and supervision from the soft labels produced by the teacher network.
 3.1 Overall Architecture
-Our network architecture, illustrated in Figure 2, comprises a 10-layer multimodal teacher GCN network and a 6-layer SNN student network.
-The teacher GCN network processes the four multimodal skeleton datasets separately.
+Our network architecture, illustrated in Figure 2, comprises a 10-layer multimodal teacher GCN modal and a 6-layer student SGN modal.
+The teacher GCN processes the four multimodal skeleton data separately.
 The processed features from each modality are then passed through the Global Average Pooling (GAP) and Fully Connected (FC) layers to generate predictions.
 These predictions are finally fused to produce the soft labels.
-The student SGN network processes each modality independently through the SSC module, transforming skeleton data into spike-form representations.
-Spike-form features are then fused via the SMF module, utilizing mutual information to optimize cross-modal integration.
-Each student network layer includes an SA-SGC module and a STC module, which are designed to capture and refine temporal and spatial features efficiently and effectively.
+The student SGN network processes each modality through the SSC module, transforming skeleton data into spike-form representations.
+Spike-form features are then fused via the SMF module, which utilizes mutual information to optimize cross-modal integration.
+Each student network layer includes the SA-SGC and STC modules, which are formulated to capture and refine spatial and temporal features efficiently and effectively.
 Our approach implements knowledge distillation through inter-layer features and soft label distillation.
-The inter-layer features from the GCN teacher network are converted into spike-form representations via the FTM, ensuring alignment with the spiking domain and facilitating the training of the SNN student network.
-Soft labels produced by the teacher network provide supplementary supervisory signals, optimizing the learning process.
-The subsequent sections provide a detailed description of the design and functionality of the modules.
+The inter-layer features from the teacher GCN modal are converted into spike-form representations via the FTM, which ensures compatibility with the spiking neural domain and facilitates the training of the student SGN modal.
+Soft labels produced by the teacher modal provide supplementary supervisory signals, optimizing the learning process.
+The subsequent sections provide a detailed account of the formulation of the modules.
 3.2 Spiking Coding Mechanism and Spiking Multimodal Fusion
-This chapter discusses how to convert continuous real skeleton data into spiking-form representation.
-Then, we describe how to effectively fuse the multimodal spike-form features derived from multimodal skeleton data.
+This chapter explores the conversion from continuous real-valued skeleton data into spiking-form representations.
+It then details the methods for fusing multimodal spike-form feature derived from multimodal skeleton data.
 (a) Multimodal Skeleton Spiking Coding and Mutual Information Weight Matrix
-(b) The detail structure of the SMIC Module
+(b) The detail structure of the Spike-form Mutual Information Calculation (SMIC) Module
 Figure 3: Multimodal Skeleton Spiking Coding and Mutual Information Weight Matrix Calculation Process
 3.2.1 Skeleton Spiking Coding Module
-For a given 3D skeleton sequence data $X_{i}\in\mathbb{R}^{C\times T\times V}$, the SSC Module first transforms it into an image-like representation.
-Then it expands it into a form suitable for spiking temporal representation illustrated in Figure 2(a).
+For a given 3D skeleton sequence data $X_{\mathrm{i}}\in\mathbb{R}^{C\times T\times V}$, the SSC Module first transforms it into an image-like representation.
+Then, it expands it into a form suitable for spiking temporal representation illustrated in Figure 2(a).
 The process is described as follows:
-$P_{0}=\operatorname{Expand}(X_{i})\in\mathbb{R}^{S\times C\times V\times T},$ (3)
-$P_{i}=\operatorname{SN}(\operatorname{BN}(\operatorname{Conv2d}(P_{0})))\in\mathbb{R}^{S\times N\times V\times T}.$ (4)
+$P_{0}=\operatorname{Expand}(X_{\text{i}})\in\mathbb{R}^{S\times C\times V\times T},$ (3)
+$P_{\text{i}}=\operatorname{SN}(\operatorname{BN}(\operatorname{Conv2d}(P_{0})))\in\mathbb{R}^{S\times D\times V\times T}.$ (4)
 In the initial step, the skeleton sequence data is transformed into an image-like representation to facilitate the structured feature extraction of spatial and temporal characteristics in subsequent stages.
-Subsequently, the $\operatorname{Expand}(\cdot)$ operation adds a spike time step dimension to the data, reorganizing the temporal structure to align with the requirements of spiking representation.
-The resulting expanded representation $P_{0}$ incorporates the additional spike time step dimension $S$ while preserving the original temporal dimension $T$, the graph structure $V$, and the channel size $C$.
-[…]
+Subsequently, the $\operatorname{Expand}(\cdot)$ operation adds a spike time step dimension $S$ to the data, reorganizing the temporal structure to align with the requirements of spiking representation.
+The expanded representation $P_{0}$ incorporates the additional spike time step dimension $S$ while preserving the original temporal dimension $T$, the graph structure $V$, and the channel size $C$.
+Following expansion, the representation $P_{0}$ is processed through a 2D convolutional layer ($\operatorname{Conv2d}$) with a kernel size of 3, followed by Batch Normalization ($\operatorname{BN}$) and a Spiking Neuron ($\operatorname{SN}$) layer.
+The pipeline refines the temporal-spatial features within the spiking domain, yielding a spiking feature representation $P$ with dimensions $S\times D\times V\times T$.
+The spike time step dimension $S$ reflects the temporal structure of spiking activity, while the dimension $T$ retains the temporal characteristics of the original sequence, ensuring a comprehensive temporal-spatial representation.
+$D$ represents the hidden channel size.
 3.2.2 Spike-based Multimodal Fusion Module
 Multimodal skeleton data consists of four components: bones ($X_{\mathrm{b}}$), joints ($X_{\mathrm{j}}$), bone motions ($X_{\mathrm{bm}}$), and joint motions ($X_{\mathrm{jm}}$), each represented with dimensions $(C,T,V)$.
-The SMF module is a mutual information-based fusion framework formulated specifically for the sparse spike-form features derived from multimodal data.
-In the SMF, the multimodal skeleton data are first encoded independently through the SSC module to obtain their respective spike-form features: $P_{\mathrm{B}},P_{\mathrm{J}},P_{\mathrm{BM}},P_{\mathrm{JM}}\in\mathbb{R}^{S\times D\times V\times T}$.
+The multimodal skeleton data are first represented independently through the SSC module to obtain their respective spike-form features: $P_{\mathrm{b}},P_{\mathrm{j}},P_{\mathrm{bm}},P_{\mathrm{jm}}\in\mathbb{R}^{S\times D\times V\times T}$ and then fed into the SMF module.
+The SMF module is based on mutual information for the fusion of sparse spike-form features derived from multimodal data.
 The spike-form features are then pairwise selected and concatenated along the feature dimension $D$ to construct a joint distribution.
 To construct a corresponding marginal distribution, one of the two spike-form features in each pair is randomly shuffled along the spike time step dimension $S$, and the same concatenation is performed.
-We propose the Spike-form Mutual Information Calculation (SMIC) network, formally constituted by an LSTM network, SN, FC, and GAP layers, to realize the spiking-driven mutual information prediction.
+We propose the Spike-form Mutual Information Calculation (SMIC) module, formally composed of LSTM, SN, FC, and GAP layers, to realize spiking-driven mutual information prediction.
 The framework computes the mutual information between four sparse spike-form features and performs weighted fusion.
-By leveraging the computational efficiency of sparse spike-form features and maximizing the lower bound of mutual information, the SMIC network effectively captures shared information across modalities, facilitating both robust and energy-efficient multimodal feature fusion within SNN framework, as illustrated in Figure 3(b).
+By leveraging the computational efficiency of sparse spike-form features and maximizing the lower bound of mutual information, the SMIC network effectively captures shared information across modalities, facilitating robust and energy-efficient multimodal feature fusion within SNN framework, as illustrated in Figure 3(b).
 To illustrate the process, we take bone ($P_{\mathrm{b}}$) and joint ($P_{\mathrm{j}}$) data as an example.
@@
 Both $P_{\mathrm{Joint}}$ and $P_{\text{marginal}}$ are then processed through an LSTM network equipped with SN layers.
-For the joint distribution, the output of the LSTM is passed through a FC layer followed by a GAP operation to compute:
+For the joint distribution, the output of the LSTM is passed through a FC layer followed by a GAP operation as:
 $\boldsymbol{t}_{1}=\operatorname{SN}(\operatorname{LSTM}(P_{\mathrm{Joint}})),\quad\boldsymbol{t}=\operatorname{GAP}(\operatorname{FC}(\boldsymbol{t}_{1})).$ (7)
@@
 $\boldsymbol{et}_{1}=\operatorname{SN}(\operatorname{LSTM}(P_{\text{marginal}})),\quad\boldsymbol{et}=\exp\left(\operatorname{GAP}(\operatorname{FC}(\boldsymbol{et}_{1}))\right).$ (8)
-Using the predictions $\boldsymbol{t}$ and $\boldsymbol{et}$, the lower bound of mutual information is estimated as:
-$mi_{-}lb=\bar{\boldsymbol{t}}-\log\overline{\boldsymbol{et}}.$ (9)
+Based on the predictions $\boldsymbol{t}$ and $\boldsymbol{et}$, the lower bound of mutual information is estimated as:
+$mi\_lb_{\mathrm{jb}}=\bar{\boldsymbol{t}}-\log\overline{\boldsymbol{et}}.$ (9)
 The calculation is performed pairwise for all modalities—bones, joints, bone motions, and joint motions—resulting in six pairwise mutual information values, which are organized into a symmetric matrix $\boldsymbol{MI}$ with zeros along the diagonal:
@@
 To normalize the mutual information values, the sum along the rows of $\boldsymbol{MI}$ is computed, yielding the average mutual information weights:
-$\boldsymbol{\overline{MI}}=\frac{\sum_{i=1}^{4}\boldsymbol{MI}_{i,:}}{\sum_{i=1}^{4}\sum_{j=1}^{4}\boldsymbol{MI}_{i,j}},$ (11)
-$\boldsymbol{W}=\frac{\boldsymbol{\overline{MI}}-\min(\boldsymbol{\overline{MI}})}{\max(\boldsymbol{\overline{MI}})-\min(\boldsymbol{\overline{MI}})}\in\mathbb{R}^{4}.$ (12)
-The fused spike-form features are then obtained by weighting the original features of all modalities using $\boldsymbol{W}$:
-$P_{\text{fuse}}=\boldsymbol{W}*\left[P_{\mathrm{B}},P_{\mathrm{J}},P_{\mathrm{BM}},P_{\mathrm{JM}}\right]^{\boldsymbol{T}}\in\mathbb{R}^{S\times D\times V\times T}.$ (13)
-Finally, the fused feature $P_{\text{fuse}}$ is fed into SGN Block for subsequent processing.
-By leveraging spiking-driven mutual information, the SMF module ensures the effective integration of sparse multimodal features, enhancing the robustness and accuracy of the overall action recognition framework.
+$\boldsymbol{\overline{MI}}=\frac{\sum_{r=1}^{4}\boldsymbol{MI}_{r,:}}{\sum_{r=1}^{4}\sum_{c=1}^{4}\boldsymbol{MI}_{r,c}},$ (11)
+$\boldsymbol{W}_{\text{MI}}=\frac{\boldsymbol{\overline{MI}}-\min(\boldsymbol{\overline{MI}})}{\max(\boldsymbol{\overline{MI}})-\min(\boldsymbol{\overline{MI}})}\in\mathbb{R}^{4}.$ (12)
+The fused spike-form features are then obtained by weighting the original features of all modalities using $\boldsymbol{W}_{\text{MI}}$:
+$P_{\text{fuse}}=\boldsymbol{W}_{\text{MI}}*\left[P_{\mathrm{b}},P_{\mathrm{j}},P_{\mathrm{bm}},P_{\mathrm{jm}}\right]^{\boldsymbol{T}}\in\mathbb{R}^{S\times D\times V\times T}.$ (13)
+Finally, the fused feature $P_{\text{fuse}}$ is fed into the SGN Block for subsequent processing.
+By leveraging spiking-driven mutual information, the SMF module ensures the effective integration of sparse multimodal features, enhancing the robustness and accuracy of the overall framework.
 3.3 SGN Block
-Revisiting vanilla skeleton-based action recognition based on GCN, we construct our SGN using features converted into spike-form and adjacency matrices. The process can be summarized as follows:
-$S=\widetilde{D}^{-\frac{1}{2}}\widetilde{A}\widetilde{D}^{-\frac{1}{2}},\quad$ (14)
-where $S$ is the normalized adjacency matrix, $\widetilde{A}=A+I$ denotes the sum of the adjacency matrix $A$ and the identity matrix $I$ (representing self-connections).
-$\widetilde{D}$ is the diagonal degree matrix of $\widetilde{A}$.
-The graph convolutional output is then calculated using the following formula:
-$H=S^{K}P\in\mathbb{R}^{S\times N\times T\times V},$ (15)
-where, $H$ represents the output of the early spatial spiking graph convolution, $K$ is the number of graph convolutional layers, and $P$ is the spike-form feature input.
-The first layer of SGC can be expressed as:
-$H^{(0)}=\mathrm{SN}(P)+\mathrm{SN}(HW_{sgc}^{(0)}),\quad$ (16)
-where $W_{sgc}^{(0)}\in\mathbb{R}^{N^{(0)}\times N^{(1)}}$ is the learnable weight matrix of the first layer.
-Each subsequent layer has its learnable weights, denoted as $W_{sgc}^{(l)}\in\mathbb{R}^{N^{(l)}\times N^{(l+1)}}$.
-To further enhance the model’s representation capabilities, we incorporate Self Spiking Attention (SSA) following the principles of existing spiking attention methods.
+Revisiting vanilla skeleton-based action recognition based on GCN, we construct our SGN with spike-form features and adjacency matrices. The process can be summarized as follows:
+$\mathcal{A}=\Delta^{-\frac{1}{2}}\widetilde{A}\Delta^{-\frac{1}{2}},\quad$ (14)
+where $\mathcal{A}$ is the normalized adjacency matrix of dimension $\times V\times V.
+KK$ represents the number of branches, each capturing distinct relationships between nodes, and $V$ denotes the number of nodes in the graph.
+$\widetilde{A}=A+I$ denotes the sum of the adjacency matrix $A$ and the identity matrix $I$ (representing self-connections), while $\Delta$ is the diagonal degree matrix of $\widetilde{A}$.
+To calculate the multi-branch graph convolution, the branch-specific adjacency matrix $\mathcal{A}_{k}\in\mathbb{R}^{V\times V}$ is applied to the input spiking-form feature $P\in\mathbb{R}^{S\times D\times T\times V}$.
+The first layer of the SGC begins by aggregating outputs from all $K$ branches and is defined as:
+$H^{(0)}=\mathrm{SN}\left(\mathrm{BN}\left(PW_{\text{r}}^{(0)}\right)\right)+\mathrm{SN}\left(\mathrm{BN}\left(\sum_{k=1}^{K}\mathcal{A}_{k}PW_{k}^{(0)}\right)\right),$ (15)
+where $\mathcal{A}_{k}$ represents the adjacency matrix for branch $k$, $W_{k}^{(0)}\text{ and }W_{\text{r}}\in\mathbb{R}^{D\times D^{\prime}}$ are the learnable weight matrix specific to the $k$-th branch and a residual connection, and $\mathrm{SN}$ applied to the aggregated features.
+To further enhance the model’s representation capabilities, we incorporate SSA following the principles of existing spiking attention methods[36].
 These methods leverage a spiking-based $QKV$ structure with scaling factors $s$ to control large matrix multiplication values.
 The computation of SSA can be described as follows:
-$Q^{(0)}=\mathrm{SN}_{Q}\left(\mathrm{BN}\left(H_{0}W^{(0)}_{Q}\right)\right),\quad$ (17)
-$K^{(0)}=\mathrm{SN}_{K}\left(\mathrm{BN}\left(H_{0}W^{(0)}_{K}\right)\right),\quad$ (18)
-$V^{(0)}=\mathrm{SN}_{V}\left(\mathrm{BN}\left(H_{0}W^{(0)}_{V}\right)\right),\quad$ (19)
-$H^{(0)}_{SA}=H_{0}+\operatorname{SN}\left((Q^{(0)}{K^{(0)}}^{T})V^{(0)}\cdot s\right),\quad$ (20)
-where, $Q^{(0)}$, $K^{(0)}$, and $V^{(0)}$ are computed through spiking neurons ($\mathrm{SN}$) after Batch Normalization ($\mathrm{BN}$) and linear transformations using learnable weights $W^{(0)}_{Q},W^{(0)}_{K},$ and $W^{(0)}_{V}$.
+$Q^{(0)}=\mathrm{SN}\left(\mathrm{BN}\left(H^{(0)}W^{(0)}_{Q}\right)\right)$$\quad$ (16)
+$K^{(0)}=\mathrm{SN}\left(\mathrm{BN}\left(H^{(0)}W^{(0)}_{K}\right)\right)$$\quad$ (17)
+$V^{(0)}=\mathrm{SN}\left(\mathrm{BN}\left(H^{(0)}W^{(0)}_{V}\right)\right)$$\quad$ (18)
+$H^{(0)}_{\text{SA}}=H^{(0)}+\operatorname{SN}\left((Q^{(0)}{K^{(0)}}^{T})V^{(0)}\cdot s\right),\quad$ (19)
+where, $Q^{(0)}$, $K^{(0)}$, and $V^{(0)}$ are computed through $\mathrm{SN}$ after $\mathrm{BN}$ and linear transformations using learnable weights $W^{(0)}_{Q},W^{(0)}_{K}$, and $W^{(0)}_{V}$.
 The attention mechanism integrates the weighted interactions, while the scaling factor $s$ prevents large values in the spiking attention computation.
-$H^{(0)}_{SA}$ combines the residual connection $H_{0}$ with the attention-enhanced feature, defined as SA-SGC, illustrated in Figure 2(b), yielding a robust representation for spiking graph processing.
-After that, the first layer of STC can be written as:
-$T^{(0)}=\operatorname{SN}\left(\mathrm{BN}\left(\operatorname{Conv2d}\left(H^{(0)}_{SA}\right)\right)\right)+\operatorname{SN}\left(H^{(0)}_{SA}\right).$ (21)
-For subsequent layers, the spiking self attention-enhanced feature $H^{(l)}_{SA}$ is updated and combined with temporal features $T^{l-1}$, expressed as,illustrated in Figure 2(c):
-$H^{(l)}_{SA}=\operatorname{SN}\left(T^{(l-1)}\right)+\operatorname{SN}\left(H^{(l-1)}_{SA}W_{sgc}^{(l-1)}\right),$ (22)
-$T^{(l)}=\operatorname{SN}\left(\mathrm{BN}^{(l)}\left(\operatorname{Conv2d}^{(l)}\left(H^{(l)}_{SA}\right)\right)\right)+\operatorname{SN}\left(H^{(l)}_{SA}\right).$ (23)
-Finally, the processed features are sent through GAP and a FC layer to output the prediction, as follows:
+$H^{(0)}_{\text{SA}}$ combines the residual connection $H^{(0)}$ with the attention-enhanced feature, defined as SA-SGC, illustrated in Figure 2(b).
+SA-SGC yields a robust representation for spiking graph processing.
+After that, the first layer of STC, illustrated in Figure 2(c), can be written as:
+$T^{(0)}=\operatorname{SN}\left(\mathrm{BN}\left(W_{\text{t}}^{(0)}\left(H^{(0)}_{\text{SA}}\right)\right)\right)+\operatorname{SN}\left(H^{(0)}_{\text{SA}}\right)$, (20)
+where $W_{\text{t}}$ is the learnable weight matrix. For subsequent layers, the spiking self attention-enhanced feature $H^{(l)}_{\text{SA}}$ and temporal features $T^{(l-1)}$, expressed as:
+$H^{(l)}=\mathrm{SN}\left(\mathrm{BN}\left(T^{(l-1)}W_{\text{r}}^{(l)}\right)\right)+\mathrm{SN}\left(\mathrm{BN}\left(\sum_{k=1}^{K}\mathcal{A}_{k}T^{(l-1)}W_{k}^{(l)}\right)\right),$ (21)
+$H^{(l)}_{\text{SA}}=H^{(l)}+\operatorname{SN}\left((Q^{(l)}{K^{(l)}}^{T})V^{(l)}\cdot s\right),\quad$ (22)
+$T^{(l)}=\operatorname{SN}\left(\mathrm{BN}^{(l)}\left(W_{\text{t}}^{(l)}\left(H^{(l)}_{\text{SA}}\right)\right)\right)+\operatorname{SN}\left(H^{(l)}_{\text{SA}}\right)$. (23)
+Finally, the output features of the ast layer $L$ are processed through GAP and a FC layer to generate the prediction, as follows:
 $y=\operatorname{FC}\left(\mathrm{GAP}\left(\operatorname{SN}\left(T^{(L)}\right)\right)\right),$ (24)
-where $y\in\mathbb{R}^{U}$, and $U$ is the number of action classes. Based on this, a standard cross-entropy loss is adopted for model optimization, formulated as:
-$L_{\mathrm{task}}=-\frac{1}{B}\sum_{b=1}^{B}\sum_{u=1}^{U}\mathrm{Y}_{bu}\log\mathrm{y}_{bu}+\left(1-\mathrm{Y}_{bu}\right)\log\left(1-\mathrm{y}_{bu}\right),$ (25)
-where, $\mathrm{Y}$ is the ground-truth label, $\mathrm{y}$ is the predicted output, and $B$ denotes the number of training samples.
+where $y\in\mathbb{R}^{U}$, and $U$ is the number of action classes.
+We further calculate the standard cross-entropy loss $\mathcal{L}_{\mathrm{task}}$ based on the predicted probabilities and one-hot encoded ground-truth labels.
 3.4 Knowledge Distillation of GCN-to-SGN
@@
 By aligning the intermediate features and outputs of the SGN with those of the GCN, shown in Figure 4(a), the method significantly enhances SGN accuracy.
-The enhancement is achieved without introducing additional computational complexity or energy consumption, thereby preserving the inherent efficiency advantages of the SGN architecture.
+The enhancement is achieved without introducing additional computational complexity or energy consumption during the inference phase, thereby preserving the inherent efficiency advantages of the SGN architecture.
 The distillation framework bridges the performance gap between GCNs and SGNs while maintaining the suitability of SGNs for energy-efficient applications.
 3.4.1 Soft Label Distillation
-In the proposed distillation framework, the CTR-GCN with a 4-stream architecture is utilized as the teacher network.
+In the proposed distillation framework, the CTR-GCN with a 4-stream architecture is utilized as the teacher network [68].
 Multimodal data inputs, including bone, joint, bone-motion, and joint-motion data, are processed independently by single-stream CTR-GCN.
-Single-modal soft labels, represented as $y_{b}$, $y_{j}$, $y_{bm}$, and $y_{jm}$, are generated corresponding to the respective modalities.
-The outputs of the four modalities are then aggregated to form the multimodal soft labels $\hat{y}$, which serve as the knowledge to be distilled into the SGN.
+Single-modal soft labels, represented as $y_{\text{b}}$, $y_{\text{j}}$, $y_{\text{bm}}$, and $y_{\text{jm}}$, are generated corresponding to the respective modalities.
+The outputs of the four modalities are then aggregated to form the multimodal soft labels $y_{\text{mm}}$, which serve as the knowledge to be distilled into the SGN.
 This aggregation is performed using weighted summation, as described below:
-$\hat{y}=\alpha_{b}y_{b}+\alpha_{j}y_{j}+\alpha_{bm}y_{bm}+\alpha_{jm}y_{jm},$ (26)
-where $\alpha_{b}$, $\alpha_{j}$, $\alpha_{bm}$, and $\alpha_{jm}$ are hyperparameters that control the relative contribution of each modality.
+$y_{\text{mm}}=\alpha_{\text{b}}y_{\text{b}}+\alpha_{\text{j}}y_{\text{j}}+\alpha_{\text{bm}}y_{\text{bm}}+\alpha_{\text{jm}}y_{\text{jm}},$ (25)
+where $\alpha_{\text{b}}$, $\alpha_{\text{j}}$, $\alpha_{\text{bm}}$, and $\alpha_{\text{jm}}$ are hyperparameters that control the relative contribution of each modality.
 These parameters are set based on the importance or reliability of the modalities during training.
-The SGN then learns to approximate the multimodal teacher soft labels $\hat{y}$ through a regression-based loss function. The final learning objective for soft label distillation is expressed as:
-$L_{sdk}=\|y-\hat{y}\|,$ (27)
+The SGN then learns to approximate the multimodal teacher soft labels $y_{\text{mm}}$ through a regression-based loss function.
+The final learning objective for soft label distillation is expressed as:
+$\mathcal{L}_{\text{sdk}}=\|y-y_{\text{mm}}\|,$ (26)
 where $y$ represents the predictions of the SGN.
@@
 To address this discrepancy, we propose the FTM, as illustrated in Figure 4(b).
-The FTM is an learnable module that transforms the continuous intermediate features of GCN into spike-form features, making them compatible with SGN.
-The FTM concatenates multimodal features along the channel dimension and fuses them using depth-wise separable convolution.
-Subsequently, continuous features are expanded.
-Finally, spike-form features are generated through Conv2D, BN, and SN layers, ensuring compatibility between the teacher GCN and student SGN architectures while preserving the valuable structural and semantic information embedded in the GCN features.
-In our distillation framework, the teacher GCN is composed of 10 GC-TC (Graph Convolution-Temporal Convolution) units, while the student SGN consists of 10 SGC-STC (Spiking Graph Convolution-Spiking Temporal Convolution) units.
-For feature distillation, we leverage intermediate-level features ($T_{\mathrm{Ab5}},T_{\mathrm{Aj5}},T_{\mathrm{Abm5}},T_{\mathrm{Ajm5}}$) and high-level features ($T_{\mathrm{Ab8}},T_{\mathrm{Aj8}},T_{\mathrm{Abm8}},T_{\mathrm{Ajm8}}$).
-The intermediate features are represented as: $T_{\mathrm{Ab5}},T_{\mathrm{Aj5}},T_{\mathrm{Abm5}},T_{\mathrm{Ajm5}}\in\mathbb{R}^{2C\times\frac{T}{2}\times V},$ while the high-level features are:$T_{\mathrm{Ab8}},T_{\mathrm{Aj8}},T_{\mathrm{Abm8}},T_{\mathrm{Ajm8}}\in\mathbb{R}^{4C\times\frac{T}{4}\times V}.$ These layers are selected because the expansion of channel dimensions and the changes in stride within the TCN network allow them to capture diverse temporal relationships and richer feature information.
-Taking the intermediate 5th layer as an example, the multimodal teacher features are first concatenated and fused as follows:
+The FTM is a learnable module that transforms the continuous intermediate features of GCN into spike-form features, making them compatible with SGN.
+[…]
 $T_{\mathrm{Am5}}=\operatorname{BN}\left(\operatorname{Conv2D}\left(\operatorname{cat}\left(T_{\mathrm{Ab5}},T_{\mathrm{Aj5}},T_{\mathrm{Abm5}},T_{\mathrm{Ajm5}}\right)\right)\right),$ (27)
@@
 To align the student SGN features with the teacher GCN features, we use cosine similarity as the loss function.
-For the intermediate 5th-layer spike-form feature and the corresponding SGN 3rd-layer spike-form feature ($S_{3}$), the loss is defined as:
-$L_{\mathrm{fkd}_{1}}=\frac{1}{J}\sum_{j=1}^{J}\left(1-\frac{T_{\mathrm{Sm5}}^{j}\cdot S_{3}^{j}}{\|T_{\mathrm{Sm5}}^{j}\|\cdot\|S_{3}^{j}\|}\right),$ (30)
-where $J$ is the number of samples. Similarly, for the high-level 8th-layer spike-form feature and the SGN 5th-layer feature ($S_{5}$):
-$L_{\mathrm{fkd}_{2}}=\frac{1}{J}\sum_{j=1}^{J}\left(1-\frac{T_{\mathrm{Sm8}}^{j}\cdot S_{5}^{j}}{\|T_{\mathrm{Sm8}}^{j}\|\cdot\|S_{5}^{j}\|}\right).$ (31)
+For the converted spike-form feature from the multimodal GCN and the corresponding 3rd-layer spike-form feature in SGN ($T^{(3)}$), the loss is defined as:
+$\mathcal{L}_{\mathrm{fkd}_{1}}=\frac{1}{J}\sum_{j=1}^{J}\left(1-\frac{T_{\mathrm{Sm5}}^{j}\cdot{T^{(3)}}^{j}}{\|T_{\mathrm{Sm5}}^{j}\|\cdot\|{T^{(3)}}^{j}\|}\right),$ (29)
+where $J$ is the number of samples. Similarly, for the high-level 8th-layer spike-form feature and the SGN 5th-layer feature ($T^{(5)}$):
+$\mathcal{L}_{\mathrm{fkd}_{2}}=\frac{1}{J}\sum_{j=1}^{J}\left(1-\frac{T_{\mathrm{Sm8}}^{j}\cdot{T^{(5)}}^{j}}{\|T_{\mathrm{Sm8}}^{j}\|\cdot\|{T^{(5)}}^{j}\|}\right).$ (30)
 The inner-layer feature distillation loss is then defined as:
-$L_{\mathrm{fkd}}=\beta_{1}L_{\mathrm{fkd}_{1}}+\beta_{2}L_{\mathrm{fkd}_{2}},$ (32)
+$\mathcal{L}_{\mathrm{fkd}}=\beta_{1}\mathcal{L}_{\mathrm{fkd}_{1}}+\beta_{2}\mathcal{L}_{\mathrm{fkd}_{2}},$ (31)
 where $\beta_{1}$ and $\beta_{2}$ are hyperparameters that control the weighting of the intermediate and high-level distillation losses.
 This formulation ensures that the student network effectively learns from the teacher’s intermediate and high-level spike-form features while optimizing the FTM for feature translation.
-Finally, the overall learning objective combines the inner-layer feature distillation loss with the task loss $L_{t}ask$ and the soft label distillation loss $L_{\mathrm{sdk}}$ as follows:
-$L=\gamma_{1}L_{\mathrm{b}}+\gamma_{2}L_{\mathrm{sdk}}+\gamma_{3}L_{\mathrm{fkd}},$ (33)
+Finally, the total loss function combines the inner-layer feature distillation loss with the task loss $\mathcal{L}_{\mathrm{task}}$ and the soft label distillation loss $\mathcal{L}_{\mathrm{sdk}}$ as follows:
+$\mathcal{L}=\gamma_{1}\mathcal{L}_{\mathrm{task}}+\gamma_{2}\mathcal{L}_{\mathrm{sdk}}+\gamma_{3}\mathcal{L}_{\mathrm{fkd}},$ (32)
 where $\gamma_{1}$, $\gamma_{2}$, and $\gamma_{3}$ are hyperparameters that balance the contribution of each loss term.
@@
 4 Experiments
-In this chapter, we present the implementation details, ablation studies, result comparisons, and visualizations, respectively. The details will be described in the following subsections.
+This chapter presents the implementation details, ablation studies, result comparisons, and visualizations. The details will be described in the following subsections.
 4.1 Experimental Setup
@@
 These datasets provide diverse action classes, extensive skeleton data, and varying levels of complexity, making them ideal benchmarks for comprehensive evaluation.
-Table 1: The details of the datasets used in our experiments.
-Dataset # Classes Split # Joints
-Cross-Subject Cross-View (CSet)
-# Train # Test # Train # Test
-NTU RGB+D 60 40,320 16,560 37,920 18,960 25
-NTU RGB+D 120 120 63,026 53,370 54,471 61,925 25
-NW-UCLA 10 - - 1,039 455 20
-NTU-RGB+D is a large-scale dataset comprising 60 action classes and 56,000 video clips.
-It includes detailed skeleton data covering a wide range of daily activities, captured using a Kinect v2 sensor.
+NTU-RGB+D [61] is a large-scale dataset comprising 60 action classes and 56,000 video clips.
+It includes detailed skeleton data covering various daily activities captured using a Kinect v2 sensor.
 The dataset provides 3D coordinates of 25 major body joints, enabling precise modeling of human movements.
 NTU-RGB+D is widely recognized as a standard benchmark in the action recognition field due to its variety of actions and large-scale data volume.
-The dataset is evaluated under two standard settings: Cross-Subject (CSub), where training and testing samples are divided based on the subjects performing the actions, and Cross-View (CView), where training and testing samples are divided based on the camera views.
-NTU-RGB+D 120, an extended version of NTU-RGB+D, significantly expands the dataset to include 120 action classes and over 114,000 video clips.
+The dataset is evaluated under two standard settings: Cross-Subject (XSub), where training and testing samples are divided based on the subjects performing the actions, and Cross-View (XView), where training and testing samples are divided based on the camera views.
+NTU-RGB+D 120 [62], an extended version of NTU-RGB+D, significantly expands the dataset to include 120 action classes and over 114,000 video clips.
 This dataset offers richer skeleton data, making it one of the largest and most comprehensive resources for skeleton-based action recognition.
 Its inclusion of a broader range of actions and more extensive data volume provides an ideal foundation for developing and evaluating advanced action recognition algorithms.
-Similarly, this dataset is evaluated under two standard settings: Cross-Subject (CSub) and Cross-Setup (CSet), where training and testing samples are divided based on the camera setups used during data collection.
-NW-UCLA is a smaller dataset consisting of 10 action classes and 1,494 video clips captured from multiple viewpoints.
+Similarly, this dataset is evaluated under two standard settings: Cross-Subject (XSub) and Cross-Setup (XSet), where training and testing samples are divided based on the camera setups used during data collection.
+NW-UCLA [63] is a smaller dataset comprising 10 action classes and 1,494 video clips captured from multiple viewpoints.
 It provides 3D coordinates of 20 major body joints, offering diverse skeleton data for assessing the robustness of action recognition models to viewpoint variations.
@@
 The specific statistics for the three datasets after the splits are summarized in Table 1.
+Table 1: The details of the datasets used in our experiments.
+Dataset # Classes Split # Joints
+Cross-Subject Cross-View (XSet)
+# Train # Test # Train # Test
+NTU RGB+D[61] 60 40,320 16,560 37,920 18,960 25
+NTU RGB+D 120[62] 120 63,026 53,370 54,471 61,925 25
+NW-UCLA[63] 10 - - 1,039 455 20
 4.1.2 Implementation Details
-The experiments are conducted on a server equipped with an NVIDIA SMI V100 GPU with 30 GB of memory.
+The experiments are conducted on a server equipped with an NVIDIA SMI V100 GPU with 32 GB of memory.
 The models are implemented using PyTorch[65] and SpikingJelly[64], which provide efficient frameworks for spiking neural networks.
-During training, we use two separate SGD optimizers: one dedicated to optimizing the SGN model and the other specifically designed to refine the FTM.
-This dual-optimizer strategy enhances the efficiency of the SGN and ensures effective feature transfer from the teacher to the student model.
 The key hyperparameters for the LIF model used in all experiments are listed in Table 2.
@@
 Surrogate function’s window size $a$ 1.0
-The training configurations for MK-SGN across different datasets are summarized in Table 3.
-We adopt an initial learning rate of 0.1, a weight decay rate of 1e-4, and set the batch size based on the dataset: 64 for NTU-RGB+D 60, 32 for NTU-RGB+D 120, and 128 for NW-UCLA.
-Training is conducted for 64 epochs with a step decay schedule, reducing the learning rate after 50 epochs.
 Table 3: Hyperparameter settings for MK-SGN training.
 Parameter NTU-RGB+D NTU-RGB+D NW-UCLA
-60 120
+120
 Learning Rate 0.1/0.1 0.1/0.1 0.1/0.1
 Decay Rate 0.1/0.1 0.1/0.1 0.1/0.1
-Batch Size 64 32 128
+Batch Size 64 64 128
 Time Steps 4 4 4
-Training Epochs 64 64 64
+Training Epochs 60 60 60
 Step 50 50 50
@@
 Weight Decay 1e-4 1e-4 1e-4
-Optimizer SGD/SGD SGD/SGD SGD/SGD
+Optimizer SGD SGD SGD
 Table 4: Hyperparameters for soft label aggregation and loss weighting.
@@
 $\gamma_{3}$ Weight of feature distillation loss term. 1.0
-In the testing phase, only the SGN model is evaluated, isolating it from the dual-optimizer framework used during training.
+The training configurations for MK-SGN across different datasets are summarized in Table 3.
+We adopt an initial learning rate of 0.1, a weight decay rate of 1e-4, and set the batch size based on the dataset: 64 for NTU-RGB+D, 64 for NTU-RGB+D 120, and 128 for NW-UCLA.
+Training is conducted for 64 epochs with a step decay schedule, reducing the learning rate to one-tenth after 50 epochs.
+In the testing phase, only the SGN model is evaluated.
 This ensures that the SGN’s performance reflects its learned capabilities without additional computational overhead.
 4.2 Ablation Studies
-To verify the effectiveness and efficiency of the proposed model, we first replace the activation layers of the classic one-stream 2S-AGCN model with LIF neurons.
+To verify the effectiveness and efficiency of the proposed model, we first replace the activation layers of the classic one-stream 2S-AGCN model [21] with LIF neurons.
 This modified model, called Base-SGN*, is then trained and tested for comparison.
-Additionally, we compare our proposed model with Spikformer, a classic SNN-based transformer architecture.
+Additionally, we compare our proposed model with Spikformer[36], a classic SNN-based transformer architecture.
 For a fair comparison, we adjust the layer structure and hidden dimension of Spikformer to match the MK-SGN architecture.
@@
 4.2.1 Theoretical Energy Consumption Calculation
-In the context of SNNs, synaptic operations (SOPs) play a critical role in determining theoretical energy consumption.
-As one of the key components of our experimental evaluation, we focus on calculating SOPs and the energy consumption of both Base-SGN* and MK-SGN, highlighting the efficiency of the proposed model.
+In the context of SNNs, synaptic operations ($\operatorname{SOPs}$) refer to the fundamental operations required to transmit spikes across synapses and perform computations within spiking neural networks.
+Specifically, $\operatorname{SOPs}$ are analogous to $\operatorname{FLOPs}$ in ANNs, serving as the core unit of the computational workload.
+Each $\operatorname{SOP}$ corresponds to a spike-based accumulate ($\operatorname{AC}$) operation, which involves multiplying a binary spike (0 or 1) by a synaptic weight and accumulating the result in the post-synaptic neuron.
+This operation leverages the simplicity of binary spike-based computations, significantly reducing computational complexity compared to ANNs.
+In ANNs, operations are defined as Multiply-and-Accumulate ($\operatorname{MACs}$), where both multiplication and addition involve floating-point numbers.
+The use of $\operatorname{SOPs}$ in spiking neural networks thus highlights their efficiency, as $\operatorname{ACs}$ require fewer computational resources than $\operatorname{MACs}$.
+As one of the critical components in our experimental evaluation, we calculate the SOPs for both Base-SGN* and MK-SGN to achieve two key objectives.
+First, this analysis underscores the advantages of our model on neuromorphic hardware by demonstrating its computational efficiency.
+Second, the calculated SOPs serve as the basis for estimating theoretical energy consumption, providing evidence of the energy-saving potential of MK-SGN.
 Table 5: Evaluation on NTU-RGB+D.
@@
 Power is the average theoretical energy consumption when predicting a skeleton from the NTU-RGB+D test set, whose calculation details are shown in Equation(35), Equation(36), and Equation(37).
-Method Architecture Param Spike Time Step SMF Knowledge Distillation SOPs FLOPs Power Ntu-RGB+D 60
-(M) (G) (G) ($mJ$) CSub(%) CView(%)
-ANN (10-layer)ST-GCN 3.1 / / / / 3.48 16.01 81.5 88.3
+Method Architecture Param Spike Time Step SMF Knowledge Distillation SOPs FLOPs Power Ntu-RGB+D
+(M) (G) (G) (mJ) Xsub(%) Xview(%)
+ANN (10-layer)ST-GCN [20] 3.1 / / / / 3.48 16.01 81.5 88.3
 SNN (10-layer)Base-SGN* 2.07 4 / / 0.60 7.17 0.536 64.2 71.3
-(6-layer)SpikformerICLR 2023 4.78 4 / / 1.69 24.07 2.17 73.9 80.6
+(6-layer)Spikformer[36]ICLR 2023 4.78 4 / / 1.69 24.07 2.17 73.9 80.6
+(10-layer) MK-SGN 3.13 4 / / 0.96 10.43 0.86 72.9 79.6
+(10-layer) MK-SGN 3.13 4 / 0.98 11.34 0.878 74.3 81.2
+(10-layer) MK-SGN 3.24 4 / 1.00 11.45 0.912 74.2 80.1
+(10-layer) MK-SGN 3.24 4 1.02 11.87 0.931 78.2 85.1
 (6-layer) MK-SGN 1.89 4 / / 0.64 7.32 0.573 74.5 80.9
-(6-layer) MK-SGN 1.89 4 / 0.66 7.32 0.582 75.9 82.1
+(6-layer) MK-SGN 1.89 4 / 0.64 7.32 0.573 75.9 82.1
 (6-layer) MK-SGN 2.17 4 / 0.67 7.76 0.596 76.1 82.6
-(6-layer) MK-SGN 2.17 4 0.68 7.76 0.614 78.5 85.6
+(6-layer) MK-SGN 2.17 4 0.67 7.76 0.596 78.5 85.6
+(4-layer) MK-SGN 1.17 4 0.37 4.13 0.212 68.5 71.6
   • 1.
     Note: FLOPs are calculated using the FVCORE library, which provides precise computation of floating-point operations.
-Following [27, 28, 29], the firing rate $r$ of each LIF neuron, is defined as the firing probability per spike time step.
-Using the firing rate $r$, the synaptic operations for a given block or layer $l$ are expressed as:
+Following [36, 37, 38], the $\operatorname{SOPs}$ for a given block or layer $l$ are defined as the product of the firing rate $r$, the number of spike time steps $S$, and the ($\operatorname{FLOPs}$) of the block or layer:
 $\operatorname{SOPs}(l)=r\times S\times\operatorname{FLOPs}(l),$ (33)
-where $\operatorname{FLOPs}(l)$ refers to the floating-point operations of $l$, defined as the number of multiply-and-accumulate(MAC) operations performed in $l$ block or layer.
-And $\operatorname{SOPs}$ are the number of spike-based accumulate (AC) operations.
 The theoretical energy consumption of Base-SGN and MK-SGN is calculated as follows:
@@
 $E_{\text{MK-SGN}}$ $=n_{m}\times E_{\text{MAC}}\times\mathrm{FL}_{\mathrm{SNNConv}}^{1}$ (35)
-$\quad+E_{\text{AC}}\times\bigg{(}\frac{k\times(k-1)}{2}\mathrm{SOP}_{SMIC}+\sum_{n=n_{m}}^{N}\mathrm{SOP}_{\mathrm{SNNConv}}^{n}$
+$\quad+E_{\text{AC}}\times\bigg(\frac{k\times(k-1)}{2}\mathrm{SOP}_{\mathrm{SMIC}}+\sum_{n=n_{m}}^{N}\mathrm{SOP}_{\mathrm{SNNConv}}^{n}$
 $\quad\quad+\sum_{m=1}^{M}\mathrm{SOP}_{\mathrm{SNNFC}}^{m}+\sum_{l=1}^{L}\mathrm{SOP}_{\mathrm{SSA}}^{l}\bigg).$
@@
 These formulations demonstrate how MK-SGN balances high performance with reduced theoretical energy consumption.
-SOPs, representing synaptic operations, serve as a critical metric for evaluating the energy efficiency of spiking-driven architectures.
 4.2.2 The Effect of Different Components
-To analyze the energy efficiency of the proposed MK-SGN and evaluate the contribution of each component within the model, we use theoretical formulas to estimate the number of operations(OPs) required for classifying a single action sample.
-Specifically, we calculate the FLOPs for ANN-based models, as well as the FLOPs for SNN on traditional GPUs and the theoretical SOPs on neuromorphic hardware.
-Using these operation counts, we further compute the energy consumption per action sample during the testing phase.
-Additionally, we evaluate the classification accuracy of MK-SGN through comprehensive ablation studies conducted on the NTU-RGB+D 60 dataset.
+We use theoretical formulas to estimate the number of operations (OPs) required for classifying a single action sample to analyse the proposed MK-SGN’s energy efficiency and evaluate each component’s contribution within the model.
+Specifically, we calculate the FLOPs for ANN-based models, the FLOPs for SNN on traditional GPUs, and the theoretical SOPs on neuromorphic hardware.
+Based on these operation counts, we further compute the energy consumption per action sample during the inference phase.
+Additionally, we evaluate the classification accuracy of MK-SGN through comprehensive ablation studies conducted on the NTU-RGB+D dataset.
 The results, including energy efficiency and accuracy comparisons, are summarized in Table 5.
 Base-SGN*
-[…]
+First, we validate the effectiveness of Base-SGN* as the basis for the ablation study.
+We compare it to the vanilla ST-GCN [20], calculating the theoretical number of operations (OPs) and energy consumption.
+The results show that Base-SGN* reduces the theoretical energy consumption to 0.536mJmJ per action sample.
+However, this increases the FLOPs and decreases the accuracy to $64.2$ and $71.3$.
+This indicates that simply replacing the activation layer is insufficient to maintain high classification accuracy.
 MK-SGN
 Building on Base-SGN*, we compare its performance with the proposed MK-SGN.
-Even without knowledge distillation, the single-modality MK-SGN achieves significant accuracy improvements of 10.1% and 8.9% on XSub and XView settings, respectively, with only a minimal increase in OPs and theoretical power consumption.
-Compared to ST-GCN, although MK-SGN demonstrates a slight increase in FLOPs, this trade-off is offset by its extremely low SOPs value of 0.68, which highlights its suitability for deployment on neuromorphic hardware.
-Furthermore, MK-SGN reduces the parameter count by over 50% and decreases FLOPs and theoretical energy consumption by more than 70% compared to structurally similar models like Spikformer, while maintaining competitive computational efficiency.
+Even without knowledge distillation, the single-modality MK-SGN achieves significant accuracy improvements of 10.3% on XSub and 9.6% on XView, with only a slight increase in OPs and theoretical power consumption.
+As shown in Table 5, the 10-layer MK-SGN achieves 72.9% accuracy on XSub and 79.6% on XView, with FLOPs at 10.43 G and power consumption at 0.86 mJ.
+Compared to ST-GCN, MK-SGN offers a better balance between accuracy and computational efficiency, significantly reducing energy consumption.
+Next, we introduce the SMF module and knowledge distillation to further enhance performance.
+The 10-layer MK-SGN with SMF and knowledge distillation achieves 78.2% accuracy on XSub and 85.1% on XView, showcasing the benefits of SMF in improving accuracy while keeping SOPs at 1.02 G and power consumption at 0.931 mJ.
+Knowledge distillation further boosts accuracy with negligible impact on computational cost.
+For the 6-layer MK-SGN, which eliminates redundant layers from the 10-layer architecture, the SMF and knowledge distillation configurations lead to an accuracy of 78.5% on XSub and 85.6% on XView.
+This version achieves low power consumption of 0.596 mJ and FLOPs of 7.76 G, demonstrating that the 6-layer model is the optimal design for balancing performance and efficiency.
+The 4-layer configuration focuses on further reducing complexity by retaining only the most critical layers (such as channel upsampling layers).
+This 4-layer MK-SGN achieves lower accuracy—68.5% on XSub and 71.6% on XView—but still benefits from the SMF and knowledge distillation techniques, achieving improvements compared to the base model.
+However, the performance gain is not as substantial as in the 6-layer design, confirming that excessive layer reduction compromises the model’s representation capacity.
+In summary, the addition of SMF and knowledge distillation enhances the accuracy of the model across all configurations.
+Among the tested architectures, the 6-layer MK-SGN strikes the best balance between accuracy and computational efficiency, demonstrating its suitability for real-time applications that require both high performance and low energy consumption.
 Table 6: Performance comparison for different modality combinations using the SMF module. Accuracy is tested on NTU-RGB+D XSub
@@
 J+B+JM+BM 2.17 0.67 0.596 76.1
-Furthermore, MK-SGN achieves an accuracy improvement of 0.6% on XSub and 0.3% on XView compared to Spikformer, demonstrating its superior balance of performance and efficiency.
-These results highlight the effectiveness of MK-SGN in achieving high accuracy with significantly reduced computational and energy requirements.
+Table 7: Comparison of different fusion methods for MK-SGN. Accuracy and theoretical energy consumption are tested on NTU-RGB+D XSub.
+Fusion Method Parameters Ops Power Accuracy
+(M) (G) (mJ) (%)
+SMF 2.17 0.67 0.596 76.1
+Spiking Attention 2.24 0.71 0.613 76.0
+Concatenation (cat) 1.90 0.64 0.589 75.4
+Sum (direct addition) 1.89 0.63 0.579 75.2
+  • 1.
+    Note:cat is performed along the channel dimension with downsampling. Spiking Attention involves pairwise modality interaction followed by summation.
 SMF
-In the case of late fusion in a multi-modal classification network, where $n_{n}$ represents the number of modalities, classifying a single action sample requires $n_{n}$ independent testing runs.
-This leads to a theoretical energy consumption scaled by $n_{n}$ times.
-Taking ST-GCN as an example, its theoretical OPs will become $4\times 3.48\,\text{G}$, and the Power will increase to $4\times 16.01\,\text{mJ}$.
-To validate the effectiveness of the designed SMF module in avoiding the increased energy consumption caused by late fusion while simultaneously improving accuracy, we conduct more in-depth experiments, as shown in Table 6.
-Specifically, we incrementally integrate different modalities such as Joint (J), Joint+Bone (J+B), Joint+Bone+Joint-Motion (J+B+JM) and Joint+Bone+Joint-Motion +Bone-Motion(J+B+JM+BM)
-[…]
-[…]
-These results validate that the SMF module effectively avoids the exponential increase in energy consumption caused by late fusion while consistently improving accuracy.
-Compared to single-modality fusion, the full multimodal fusion achieves an improvement of $1.6\%$ in accuracy for CSub, showcasing the robustness of the SMF module in efficiently integrating information across multiple modalities.
-This balance of efficiency and accuracy highlights the superiority of SMF over traditional late fusion methods.
+In multi-modal classification, late fusion requires independent inference for each modality, causing energy consumption to increase by a factor of $n$ for $n$ modalities.
+For example, fusing four modalities—Joint (J), Bone (B), Joint Motion (JM), and Bone Motion (BM) in ST-GCN [20]—increases operations to $4\times 3.48\,\text{G}$ and power consumption to $4\times 16.01\,\text{mJ}$.
+To reduce this energy cost while maintaining accuracy, we introduce the SMF module, which integrates modalities more efficiently.
+Table 6 shows the results from incremental fusion of modalities: J, J+B, J+B+JM, and J+B+JM+BM.
+For the single Joint modality, the accuracy is $74.5\%$ with $0.573\,\text{mJ}$ of power consumption.
+As more modalities are added, accuracy increases to $76.1\%$ with a slight rise in energy consumption to $0.596\,\text{mJ}$.
+The number of parameters increases from $1.89\,\text{M}$ to $2.17\,\text{M}$, and operations from $0.64\,\text{G}$ to $0.67\,\text{G}$, showing that the SMF module efficiently integrates modalities without significantly increasing complexity.
+These results confirm that the SMF module avoids the linear increase in energy consumption associated with late fusion, while improving accuracy by $1.6\%$ compared to single-modality fusion.
+This demonstrates the SMF module’s effectiveness in integrating multiple modalities efficiently.
+Comparison with Other Fusion Methods
+To further assess the effectiveness of the SMF module, we compare it with simpler fusion methods, including Spiking Attention, Concatenation (cat), and direct sum (Sum), using the same four modalities: Joint (J), Bone (B), Joint Motion (JM), and Bone Motion (BM).
+These methods are commonly used in multimodal fusion tasks but often suffer from higher computational overhead.
+The comparison results are summarized in Table 7.
+Spiking Attention and Concatenation (cat) show similar accuracy improvements compared to single-modality fusion, with accuracies of $75.4\%$ and $76.0\%$, respectively.
+However, both methods require more computational resources than SMF.
+The Sum method, which directly adds the modality features, results in the lowest accuracy of $75.2\%$, as it fails to effectively model the interactions between modalities.
+The SMF module achieves the highest accuracy of $76.1\%$ with relatively low increase in power consumption ($0.596\,\text{mJ}$), making it the most efficient method in terms of both accuracy and energy consumption.
+This validates the superiority of SMF over traditional late fusion techniques, as it balances accuracy improvements with computational efficiency.
 Knowledge Distillation of GCN-to-SGN
+We evaluate the impact of our proposed knowledge distillation method by comparing SGN models trained with and without teacher guidance.
+The distillation method consists of Soft Label Distillation (SKD) and Inner-layer Feature Distillation (IFD).
+Ablation experiments are conducted to assess the contributions of each component.
+Additionally, we examine the effect of SKD using different modality combinations and the influence of IFD at various levels (low-level, high-level, or both).
+The results in Table 8 demonstrate the effectiveness of SKD and IFD.
+With SKD, accuracy improves as more modalities are incorporated: single-modality achieves 75.1%, two-modalities 75.4%, and four-modalities 75.6%, highlighting the benefit of multimodal knowledge transfer.
+For IFD, focusing on high-level features yields better performance (75.3%) than low-level features (74.9%), indicating the significance of abstract information in guiding the student model.
+Combining both SKD and IFD in the Full Knowledge Distillation (Full KD) setup achieves 75.9% accuracy, demonstrating the advantage of integrating soft-label guidance with feature-level distillation.
+These results emphasize the importance of combining output-level and feature-level distillation to enhance SGN performance, with SKD aligning the student’s outputs with the teacher’s predictions, and IFD enabling the student to better emulate the teacher’s intermediate representations.
 Table 8: Accuracy comparison under different knowledge distillation settings. Accuracy is tested on NTU-RGB+D XSub
@@
 - - - - 74.5
-Soft Label Only - J 75.1
+Soft Label (Single Modal.) - J 75.1
 Soft Label (Two Modal.) - J+B 75.4
@@
 Full KD J+B+JM+BM 75.9
-We compare the accuracy of SGN models trained with and without teacher guidance to evaluate the impact of our proposed knowledge distillation method.
-Knowledge distillation consists of two components: soft label distillation (SKD) and inner-layer feature distillation (IFD).
-To further analyze their contributions, we conduct experiments with and without each component.
-Additionally, we investigate the impact of soft label distillation using different modality combinations, as well as the effect of inner-layer feature distillation applied at different levels (low-level, high-level, or both).
-The results in Table 7 demonstrate the significant contributions of SKD and IFD to the accuracy improvements of SGN models.
-With the introduction of SKD, accuracy improves steadily as more modalities are incorporated.
-Single-modality SKD achieves 75.1%, while combining two and four modalities further enhances the performance to 75.4% and 75.6%, respectively.
-This progression highlights the importance of multimodal knowledge transfer from the teacher network to the student model.
-For IFD, focusing on low-level features yields an accuracy of 74.9%, while high-level features achieve 75.3%, indicating that the more abstract information encoded in high-level features plays a slightly greater role in guiding the student model.
-However, combining both low-level and high-level feature distillation demonstrates their complementary effects, achieving a final accuracy of 75.9% when combined with SKD in the full knowledge distillation (Full KD) setup.
-The results highlight the importance of leveraging both output-level and feature-level guidance.
-While SKD aligns the student’s outputs with the teacher’s multimodal predictions, IFD bridges the gap in intermediate representations, allowing the student to better emulate the teacher’s internal structure.
-The findings validate the proposed knowledge distillation framework as an effective means of improving SGN model performance through both comprehensive guidance and multimodal integration.
+Table 9: Sensitivity Analysis of Model Hyperparameters and Accuracy for MK-SGN.
+Model Hyperparameters Accuracy (%)
+$\alpha_{b}$ $\alpha_{j}$ $\alpha_{bm}$ $\alpha_{jm}$ $\beta_{1}$ $\beta_{2}$ $\gamma_{1}$ $\gamma_{2}$ $\gamma_{3}$
+0.1 0.1 0.1 0.1 0.5 0.5 0.5 0.5 0.5 77.1
+0.5 0.5 0.5 0.5 0.5 0.5 1.0 1.0 1.0 77.5
+0.25 0.5 0.25 0.25 0.5 0.5 0.8 0.9 1.0 78.3
+0.75 0.25 0.75 0.75 0.5 0.5 0.6 0.7 1.0 77.9
+0.25 0.25 0.5 0.5 0.5 0.5 0.9 1.0 1.0 78.0
+0.25 0.25 0.25 0.25 0.5 0.5 1.0 1.0 1.0 78.5
+0.25 0.25 0.25 0.25 1.0 1.0 0.7 0.6 0.5 77.6
+0.1 0.25 0.25 0.25 1.0 0.5 0.8 0.9 0.6 77.5
 Integrated Module Performance Comparison
 We carefully evaluate the contributions of each component in the MK-SGN model, with results summarized in Table 5.
-The fully integrated MK-SGN achieves accuracies of 78.5% and 85.6% on XSub and XView settings, respectively, with a spike time step of 4.
-Compared to Spikformer, MK-SGN reduces the parameter count by 54.6% and FLOPs by 67.4%, significantly enhancing its feasibility for training on traditional GPUs.
-Moreover, MK-SGN decreases SOPs by 59.76% and theoretical energy consumption by 71.7%, while achieving accuracy improvements of 4.6% and 5.0%, demonstrating its ability to balance high performance with low energy consumption.
+The fully integrated MK-SGN achieves accuracy of 78.5% and 85.6% on XSub and XView settings, respectively, with a spike time step of 4.
+Compared to Spikformer, MK-SGN reduces the parameter count by 54.6% and FLOPs by 67.7 %, significantly enhancing its feasibility for training on traditional GPUs.
+Moreover, MK-SGN decreases SOPs by 60.35% and theoretical energy consumption by 72.53% while achieving accuracy improvements of 4.6% and 5.0%, demonstrating its ability to balance high performance with low energy consumption.
 Compared to Base-SGN*, MK-SGN delivers substantial accuracy gains of 14.3% and 11.3% on XSub and XView, respectively.
-Additionally, when compared to the classic ANN-based model ST-GCN, MK-SGN achieves a remarkable 94.78% reduction in energy consumption while maintaining comparable accuracy.
+Additionally, when compared to the classic ANN-based model ST-GCN, MK-SGN achieves a remarkable 96.28% reduction in energy consumption while maintaining comparable accuracy.
 These results highlight the efficiency and effectiveness of MK-SGN, showcasing its suitability for energy-constrained applications without compromising performance.
-4.3 Comparison with State-of-the-Art Methods
-In this section, we conduct a comprehensive comparison between our proposed MK-SGN and SOTA methods for skeleton-based action recognition.
-The comparison is divided into two parts:(1).
-model architecture comparison and (2).metric comparison.
-In the model architecture comparison, we focus on the design and layer configurations of MK-SGN in contrast to GCN and SNN models, using the NTU-RGB+D 60 dataset as a case study.
-This includes both publicly available results and those we reproduced under consistent experimental settings.
-The architecture analysis establishes the foundation for understanding the computational and operational differences, which are crucial for fair metric comparisons.
-In the metric comparison, we evaluate MK-SGN against SOTA models across several dimensions, including accuracy, number of operations (OPs), and theoretical energy consumption.
-These metrics provide a quantitative perspective on MK-SGN’s ability to balance high performance with energy efficiency, particularly in energy-constrained environments.
-By integrating both structural and performance analyses, we aim to demonstrate the superiority of MK-SGN in achieving competitive results with significantly reduced energy and computational overhead, establishing its practicality and effectiveness in real-world applications.
-Table 8: Model’s architecture for NTU-RGB+D 60
+Sensitivity Analysis Experiments
+After evaluating the effect of different components in MK-SGN, we conduct sensitivity analysis on key model hyperparameters to assess their impact on accuracy.
+Unlike training hyperparameters, which have predictable effects, model parameters such as $\alpha$, $\beta$, and $\gamma$ directly influence model performance.
+Experiments are conducted on the NTU-RGB+D XSub dataset.
+As shown in Table 9, different configurations of the hyperparameters are tested.
+The best performance, achieving 78.5% accuracy, results from a balanced configuration ($\alpha_{b}=0.25,\alpha_{j}=0.25,\alpha_{bm}=0.25,\alpha_{jm}=0.25$, $\beta_{1}=\beta_{2}=0.25$ and $\gamma_{1}=\gamma_{2}=\gamma_{3}=1.0$).
+The results demonstrate that MK-SGN’s accuracy is sensitive to these model hyperparameters.
+Variations from the optimal configuration lead to a decrease in accuracy, highlighting the importance of tuning these parameters for optimal performance.
+4.3 Visualization Analysis
+Spiking Self-Attention Visualization
+To elucidate how MK-SGN makes decisions and to quantify the contribution of spiking attention across depth and time, we visualize the spike self-attention maps produced by the SSA module, in Fig.5.
+Concretely, we render the aggregated $k\!\odot\!v$ maps (summed over heads) at each layer and along the temporal axis; higher grayscale values indicate stronger spike-modulated attention.
+The horizontal axis indexes the NTU 25 joints, and the vertical axis indexes time steps $T$.
+(a) Spiking attention maps for Drop.
+(b) Spiking attention maps for Put on Glasses.
+(c) Spiking attention maps for Hopping.
+Figure 5: Spiking attention maps for different actions.
+The x-axis represents NTU 25 joints, and the y-axis corresponds to time steps $T$.
+The grayscale intensity indicates the activation level of the attention mechanism at each joint over time.
+Case A: Drop.
+In Fig.
+5(a), early layers (L1–L2) exhibit diffuse activations.
+From L3, attention concentrates on the upper-limb chain (shoulder–elbow–wrist–hand, joints 5–12), which directly governs the dropping motion.
+Moderate responses on the spine chain (joints 1–3, 21) indicate posture stabilization, while the lower-limb chains (joints 13–20) remain weak.
+In L6, activity condenses into sparse temporal bands aligned with release and follow-through, showing SSA’s ability to emphasize task-critical skeletal substructures.
+Case B: Put on Glasses.
+In Fig.
+5(b), initial activations are diffuse.
+From L3, attention intensifies on the hand–arm chains (shoulder–elbow–wrist–hand, joints 5–8, 9–12) and their connection to the head chain (neck–head, joints 3–4, 21).
+This highlights the coordinated interaction of arms and head required to lift and place glasses.
+In later layers, spikes are temporally localized around approach and placement, emphasizing fine-grained motor control in upper-limb–head coordination.
+Case C: Hopping.
+As shown in Fig.
+5(c), early layers (L1–L2) present distributed activations.
+From L3, focus shifts toward the lower-limb chains (hip–knee–ankle–foot, joints 13–20), which dominate propulsion and landing.
+In deeper layers, the spine chain (joints 1–3, 21) gains attention, reflecting its stabilizing role during mid-air balance.
+Final layers show temporally sharp peaks near takeoff and landing, matching the energy-intensive phases of hopping.
+This analysis reveals how the attention mechanism evolves to prioritize different joints depending on the action, with the model focusing more on the relevant joints at each phase of the action.
+Visualization of CAMs on Temporal and Spatial Dimensions
+To demonstrate the MK-SGN model’s effectiveness in capturing key features, we apply Class Activation Mapping (CAM) to the outputs of the SGC and STC layers, as shown in Fig.6.
+The spatial CAM highlights activations from the SGC layer, focusing on relevant joints, while the temporal CAM reflects activations from the STC layer, emphasizing critical time steps.
+This visualization clearly shows how the model attends to the most important body parts and action phases for accurate classification.
+Figure 6: CAM visualizations on spatial and temporal dimensions.
+Each subfigure corresponds to one action class, with the upper skeleton plot showing CAM intensities across spatial joints and the lower bar showing temporal contributions.
+Warmer colors (yellow/white) denote stronger activation responses, indicating which joints and time frames are most critical for the model’s classification.
+Case A: Drop.
+For the Drop action, the CAM in Fig.6(a) shows that MK-SGN focuses on the upper-limb joints, particularly the shoulders, elbows, and wrists, which are crucial during the release phase.
+The trunk joints are moderately activated, reflecting their role in posture stabilization, while the lower-limb joints receive slight attention.
+In the temporal domain, peak activations occur around frames 7–8, when the object is released.
+Case B: Put on Glasses.
+For the Put on Glasses action (Fig.6(b)), the CAM highlights the face-related joints, especially the eyes, which receive the highest activation, aligning with their significance in the action.
+The lower limbs and trunk show little attention, reflecting their relatively low involvement.
+In the temporal domain, activations peak around frames 7–8, capturing the critical moment when the glasses are placed.
+Case C: Hopping.
+For the Hopping action (Fig.6(c)), MK-SGN concentrates attention on the lower limbs, particularly the knees and ankles, which dominate the motion.
+The upper body receives slight attention, consistent with the nature of the action.
+In the temporal domain, intermittent peaks are observed, especially near frames 9–10, reflecting the recurrent critical moments of the take-off process captured by the model.
+In all three cases, MK-SGN’s CAM visualizations confirm its ability to focus on the most relevant spatial and temporal features for each action.
+The model selectively attends to key joints and critical time steps, proving its effectiveness in capturing the necessary cues for accurate action recognition.
+t-SNE Visualization of Action Embeddings.
+Figure 7: t-SNE visualization of action embeddings.
+The plot shows the embeddings of six action samples across four variants of MK-SGN: (a) Base-SGN, (b) MK-SGN, (c) MK-SGN (+SMF), and (d) MK-SGN (+SMF+Full KD).
+The clusters demonstrate the effectiveness of each model variant in capturing and distinguishing actions representation.
+Fig. 7 presents the t-SNE visualization of six sample actions across four MK-SGN variants: Base-SGN, MK-SGN, MK-SGN (+SMF), and MK-SGN (+SMF+Full KD).
+For Base-SGN, the embeddings show scattered distributions with slight separation between actions, indicating limited feature discrimination.
+MK-SGN, on the other hand, demonstrates significant improvement, with clearer clusters forming around individual actions, highlighting the effectiveness of spiking graph convolution and attention mechanisms.
+The addition of SMF further refines the clusters, showing improved separation between actions.
+Finally, the Full KD variant achieves the most compact clusters, suggesting that knowledge distillation enhances the model’s ability to learn more discriminative action representations.
+This analysis underscores the incremental improvements in action representation, validating the effectiveness of each component in enhancing the model’s ability to capture and distinguish complex action features.
+Table 10: Model’s architecture for NTU-RGB+D
 Layer GCN GCN Structure Base-SGN MK-SGN SGN Structure Spikformer Spikformer Structure
@@
 10 256/256 GC-TC 256/256 - SGC-STC - -
-final 256/class number GAP-FC 256/class number 256/class number GAP-SN-FC 256/class number GAP-FC
+final 256/class number GAP-FC 256/class number 256/class number SN-GAP-FC 256/class number SN-GAP-FC
+4.4 Comparison with State-of-the-Art Methods
+This section comprehensively compares our proposed MK-SGN and SOTA methods for skeleton-based action recognition.
+The comparison is divided into two parts: (1).model architecture comparison and (2).metric comparison.
+In the model architecture comparison, we focus on the configuration and layer design of MK-SGN in contrast to vanilla GCN and SNN models, using the NTU-RGB+D 60 dataset as a case study.
+By analyzing the architectural differences, this comparison provides a foundation for the subsequent evaluation of performance metrics, ensuring a fair and consistent basis for comparison.
+In the metric comparison, we evaluate MK-SGN against SOTA models across several dimensions, including accuracy, number of operations (OPs), and theoretical energy consumption.
+This evaluation encompasses publicly available results and those reproduced under consistent experimental settings, providing a comprehensive perspective on MK-SGN’s ability to balance high performance with energy efficiency, particularly in energy-constrained environments.
+By integrating both structural and performance analyses, we aim to demonstrate the superiority of MK-SGN in achieving competitive results with significantly reduced energy and computational overhead, establishing its practicality and effectiveness in real-world applications.
 4.4.1 Model Architecture Comparison
-Table 8 summarizes the architectural differences between our proposed MK-SGN and various SOTA models, including GCN, and Spikformer, as applied to the NTU-RGB+D 60 dataset.
-This comparison highlights how each model is structured to handle skeleton-based action recognition tasks, with adaptations made to ensure fair performance evaluations.
-The GCN model serves as a classical ANN-based approach, employing GC layers combined with TC layers for spatio-temporal feature extraction.
-Its structure consists of 10 layers of GC-TC operations, with progressively increasing feature dimensions and downsampling via strides, culminating in a GAP-FC layer for classification.
-This architecture represents a strong baseline for ANN-based recognition.
-The Base-SGN* model acts as the baseline SNN structure, leveraging SGC and STC layers to process spike-form data.
-It follows a similar configuration to GCN but is adapted for spiking computation.
-However, it lacks the advanced features introduced in MK-SGN.
-The Spikformer model, a spiking-based Transformer, incorporates SSA layers and MLP components for feature extraction.
+Table 10 summarizes the architectural differences between our proposed MK-SGN and various models, including the unified framework of Vanilla GCN [20, 21, 22, 23] and Spikformer [36], as applied to the NTU-RGB+D dataset.
+This comparison highlights how each model is structured to handle skeleton-based action recognition tasks.
+The Vanilla GCN model is a classic ANN-based approach, utilizing GC layers to capture spatial dependencies and TC layers for temporal modeling.
+Its architecture typically follows a pyramid design with 10 layers of GC-TC operations, featuring progressively increasing feature dimensions and temporal downsampling through strides, culminating in a GAP-FC layer for final classification.
+The Base-SGN* model serves as the baseline SNN structure, leveraging SGC and STC layers to process spike-form data. It follows a similar configuration to GCN but is adapted for spiking computation.
+The Spikformer model, a spiking-based Transformer, incorporates SSA layers and Multilayer Perceptron (MLP) components for feature extraction.
 For this study, we modified Spikformer to align with MK-SGN in terms of layer count and hidden dimensions, ensuring a fair comparison of metrics such as energy efficiency and accuracy.
 This adjustment enables us to evaluate the performance of Spikformer in a setting directly comparable to MK-SGN.
-Similarly, other network variations like Spiking-driven-former, Spiking-driven-former v2 and Spiking Wavelet Transformer used in the study are adapted to maintain consistency across evaluations.
+Similarly, other network variations, such as Spiking-driven-former, Spiking-driven-former V2, and Spiking Wavelet Transformer, used in the study are adapted to maintain consistency across evaluations.
 Finally, the MK-SGN model stands out with its innovative spiking-specific modules, including the SSC and SMF modules, which enhance multimodal integration and energy efficiency.
@@
 4.4.2 Metric Comparison
-Table 9: Comparative results on NTU RGB+D 60 and 120. We evaluate our model in terms of classification accuracy (%). X-Sub and X-View represent cross-subject and cross-view splits.
+Table 11: Comparative results on NTU RGB+D and NTU RGB+D 120.
+We evaluate our model in terms of classification accuracy (%).
+XSub, XView and XSet represent Cross-Subject, Cross-View and Cross-Setup splits.
 Method Architecture Backbone Spike Time Step Performance Accuracy (%)
-Param OPs Power Csub Cview Csub CSet
-(M) (G) (mJ) (60) (60) (120) (120)
+Param OPs Power XSub XView XSub XSet
+(M) (G) (mJ) (120) (120)
 Lie Group[67] ANN CNN / / / / 50.1 52.8 / /
@@
 ST-GCN [20] ANN GCN / 3.1 3.48 16.01 81.5 88.3 70.7 73.2
-2S-AGCN [12] ANN GCN / 3.48 35.8 164.68 88.5 95.0 82.5 84.2
-Shift-GCN [13] ANN GCN / / 10 46 90.7 96.5 85.9 87.6
-MS-G3D [59] ANN GCN / 3.19 67.63 311.09 91.5 96.2 86.9 88.4
+2S-AGCN [21] ANN GCN / 3.48 37.32 182.87 88.5 95.0 82.5 84.2
+Shift-GCN [22] ANN GCN / / / / 90.7 96.5 85.9 87.6
+MS-G3D [68] ANN GCN / 3.19 48.88 239.51 91.5 96.2 86.9 88.4
+CD-JBF-GCN[5] ANN GCN / 1.46 / / 89.0 95.7 / /
 CTR-GCN[23] ANN GCN / 1.46 7.88 36.25 92.4 96.8 88.9 90.6
+Info-GCN [24] ANN GCN / 1.6 7.36 33.86 92.3 96.7 89.2 90.7
+FR-Head[4] ANN GCN / 2.0 / / 92.8 96.8 89.5 90.9
+BlockGCN[7] ANN GCN / 1.3 7.08 32.57 93.1 97.0 90.3 91.5
 Base-SGN* SNN SGN 4 2.07 0.60 0.536 64.2 71.3 45.2 47.3
@@
 Spiking Wavelet Transformer[69]ECCV 2025 SNN Spiking Transformer 4 3.24 1.48 2.01 74.7 81.2 63.5 64.7
-MK-SGN (Ours) SNN SGN 4 2.17 0.68 0.614 78.5 85.6 67.8 69.5
+STAtten[9]CVPR 2025 SNN Spiking Transformer 4 3.19 1.98 2.48 72.8 79.7 60.3 61.7
+MK-SGN (Ours) SNN SGN 4 2.17 0.67 0.596 78.5 85.6 67.8 69.5
   • 1.
@@
 CTR-GCN [23] ANN GCN / 1.46 2.48 12.15 96.5
+Info-GCN [24] ANN GCN / 1.6 2.31 10.64 96.5
+FR-Head[4] ANN GCN / 2.0 / / 96.8
+BlockGCN[7] ANN GCN / 1.3 1.96 9.05 96.9
 Base-SGN* SNN SGN 4 2.07 0.151 0.195 86.4
@@
 Spiking Wavelet Transformer[69]ECCV 2025 SNN Spiking Transformer 4 3.24 0.451 0.697 86.9
+STAtten[9]CVPR 2025 SNN Spiking Transformer 4 3.19 0.451 0.697 86.9
 MK-SGN (Ours) SNN SGN 4 2.17 0.165 0.207 92.3
@@
     Note: FLOPs are calculated using the FVCORE library, which provides precise computation of floating-point operations.
-We compare our proposed MK-SGN with the current SOTA ANNs and SNNs in skeleton-based action recognition, focusing on two critical performance metrics: OPs and theoretical energy consumption.
+We compare our proposed MK-SGN with recent state-of-the-art (SOTA) ANNs and SNNs in skeleton-based action recognition, focusing on two key metrics: OPs and theoretical energy consumption.
 The results are summarized in Table 11 and Table 12.
-To ensure fairness, the reported results for existing multi-modal GCN methods utilize all four modalities: joint, bone, joint motion, and bone motion.
-For these methods, the softmax scores of multiple streams are fused by summing the scores from each stream.
-The OPs and theoretical energy consumption for these multi-modal methods are calculated by multiplying the values for a single stream by the total number of modalities.
-Additionally, we reproduce the results of SOTA SNN models using only the joint modality to provide a direct and fair comparison with MK-SGN.
-These reproduced SNN models are carefully adapted to align with MK-SGN in terms of the number of layers and hidden dimensions.
-This ensures consistency in the experimental settings and allows for a robust evaluation of their efficiency and effectiveness.
-Figure 5: Theoretical Energy Consumption Comparison
-(a) Comparsion on NTU-RGB+D 60 Csub
-(b) Comparsion on NTU-RGB+D 60 Cview
-(c) Comparsion on NTU-RGB+D 120 Csub
-(d) Comparsion on NTU-RGB+D 120 Cset
-Figure 6: Comparsion of Accuracy with SOTA SNN
-The comparison highlights that MK-SGN strikes a remarkable balance between accuracy and energy efficiency.
-Compared to existing methods, MK-SGN achieves significant reductions in theoretical energy consumption, emphasizing its suitability for energy-constrained scenarios, such as edge computing and real-time applications.
-By leveraging its spiking-driven architecture, MK-SGN demonstrates the ability to deliver competitive or superior performance while maintaining exceptional energy efficiency.
-Detailed results and analysis are provided in the following subsections.
-Results on NTU-RGB+D 60 and 120
-As shown in Table 9, MK-SGN demonstrates superior performance on both NTU-RGB+D 60 and NTU-RGB+D 120 datasets, showcasing its remarkable energy efficiency and competitive accuracy.
-On the NTU-RGB+D 60 and NTU-RGB+D 120 datasets, MK-SGN achieves a significant reduction in theoretical energy consumption, requiring only 0.614 mJ per sample.
-This represents a 96.16% decrease compared to ST-GCN (16.01 mJ), a 99.63% decrease compared to 2S-AGCN (164.68 mJ), and a 99.80% decrease compared to MS-G3D (311.09 mJ).
-Even when compared to lightweight models like Shift-GCN (46 mJ) and CTR-GCN (36.25 mJ), MK-SGN reduces energy consumption by 98.67% and 98.31%, respectively, highlighting its remarkable efficiency advantage across all ANN-based GCN architectures, as showed in Figure 5.
-Beyond energy savings, MK-SGN surpasses SOTA SNN-based models in both accuracy and efficiency, as highlighted in Table 9 and Figure 6.
-For NTU-RGB+D 60, MK-SGN achieves classification accuracies of 78.5% (Csub) and 85.6% (Cview), outperforming competitive SNN models such as Spikformer (73.9% Csub, 80.1% Cview) and Spike-driven Transformer V2 (77.4% Csub, 83.6% Cview).
-Importantly, MK-SGN achieves this with significantly reduced energy consumption, requiring only 0.614 mJ per sample compared to Spikformer’s 2.17 mJ and Spike-driven Transformer V2’s 2.91 mJ.
-MK-SGN also demonstrates an efficient parameter design with 2.17M parameters, far less than Spike-driven Transformer V2’s 11.47M, highlighting its lightweight and efficient architecture.
-On the NTU-RGB+D 120 dataset, MK-SGN continues to outperform SNN models, achieving accuracies of 67.8% (Csub) and 69.5% (CSet).
-In comparison, Spikformer achieves 61.7% (Csub) and 63.7% (CSet), while Spike-driven Transformer V2 achieves 64.3% (Csub) and 65.9% (CSet).
-Despite these improvements in accuracy, MK-SGN remains highly energy-efficient, consuming only 0.614 mJ, compared to Spiking Wavelet Transformer’s 2.01 mJ and Spike-driven Transformer V2’s 2.91 mJ.
-This demonstrates MK-SGN’s superior performance and efficiency across datasets.
-These results validate MK-SGN as a robust and energy-efficient solution for skeleton-based action recognition, excelling in scenarios requiring both high accuracy and low computational cost.
-By leveraging its spiking-driven architecture, MK-SGN sets a new benchmark for balancing performance and efficiency in SNN.
-Results on NW-UCLA
-On the NW-UCLA dataset, as summarized in Table 10, MK-SGN achieves a classification accuracy of 92.3%, surpassing SOTA SNN models.
-It outperforms models like Spikformer (85.4%) and Spike-driven Transformer V2 (89.4%) while maintaining significantly lower energy consumption, consuming only 0.207 mJ per sample compared to 0.914 mJ for Spike-driven Transformer V2.
-This highlights MK-SGN’s ability to effectively balance accuracy and energy efficiency without compromising its computational performance.
-Its lightweight structure and optimized spiking modules allow for efficient recognition, particularly in scenarios where minimizing resource consumption is critical.
-These results further emphasize MK-SGN’s versatility and practicality across diverse application settings.
+To ensure a fair comparison, results for existing multi-modal GCN methods use the maximum number of modalities specified in their respective papers, including joint, bone, joint motion, and bone motion.
+The OPs and theoretical energy consumption for these multi-modal methods are calculated by multiplying the single-stream values by the number of modalities used.
+Additionally, we reproduce the results of recent SOTA SNNs using only the joint modality for a direct comparison with MK-SGN.
+These models are adapted to align with MK-SGN in terms of layers and hidden dimensions, ensuring consistency across experimental settings.
+The comparison, shown in Table 11, highlights MK-SGN’s exceptional balance between accuracy and energy efficiency.
+MK-SGN significantly reduces theoretical energy consumption, requiring only 0.596 mJ per sample, representing a 96.28% reduction compared to ST-GCN (16.01 mJ), 99.67% reduction compared to 2S-AGCN (182.87 mJ), and 99.75% reduction compared to MS-G3D (239.51 mJ).
+For lightweight models like Shift-GCN (46 mJ), CTR-GCN (36.25 mJ) and BlockGCN (32.57 mJ), MK-SGN reduces energy consumption by 98.70%, 98.36% and 98.75%, respectively, demonstrating its superior efficiency across ANN-based GCN models.
+Beyond energy savings, MK-SGN outperforms recent SNN models in both accuracy and efficiency.
+For NTU-RGB+D, MK-SGN achieves 78.5% (XSub) and 85.6% (XView), surpassing competitive SNN models such as Spikformer (73.9% XSub, 80.1% XView) and Spike-driven Transformer V2 (77.4% XSub, 83.6% XView).
+Notably, MK-SGN achieves this with significantly reduced energy consumption—only 0.596 mJ per sample, compared to Spikformer’s 2.17 mJ and Spike-driven Transformer V2’s 2.91 mJ.
+Furthermore, MK-SGN demonstrates a highly efficient design, with only 2.17M parameters, far fewer than Spike-driven Transformer V2’s 11.47M.
+On the NTU-RGB+D 120 dataset, MK-SGN continues to outperform SNN models, achieving 67.8% (XSub) and 69.5% (XSet), while Spikformer and Spike-driven Transformer V2 achieve lower accuracies, 61.7% (XSub) and 63.7% (XSet), and 64.3% (XSub) and 65.9% (XSet), respectively.
+This illustrates MK-SGN’s ability to maintain competitive performance and efficiency across different datasets.
+On the NW-UCLA dataset, as shown in Table 12, MK-SGN achieves 92.3% accuracy, surpassing other SOTA SNN models like Spikformer (85.4%) and Spike-driven Transformer V2 (89.4%), while maintaining significantly lower energy consumption (0.207 mJ per sample), compared to Spike-driven Transformer V2 (0.914 mJ).
+This confirms MK-SGN’s remarkable ability to balance accuracy and energy efficiency, especially in scenarios that demand low computational cost.
+To further illustrate the synergy between accuracy and energy efficiency, we present the model comparison in Figure 8.
+The results, based on the NTU-RGB+D CS dataset, highlight MK-SGN’s superior performance in terms of both accuracy and power consumption.
+These results validate MK-SGN as a reliable and energy-efficient solution for skeleton-based action recognition, excelling in scenarios requiring high accuracy and low computational overhead.
+The spiking-driven architecture of MK-SGN sets a new benchmark for balancing performance and efficiency in SNNs.
+Figure 8: Comparison of accuracy and power consumption for various models on the NTU-RGB+D CS dataset. MK-SGN outperforms other models in both accuracy and energy efficiency.
+4.5 Discussion
+Principle Analysis: Accuracy and Efficiency Trade-off
+MK-SGN achieves significant energy efficiency, but its accuracy is limited by the sparsity of SNNs.
+Unlike ANNs, which transmit information continuously, SNNs only transmit during spikes, leading to sparse information transfer and lower entropy.
+Information entropy $H(X)$ is defined as:
+$H(X)=-\sum_{i=1}^{n}p_{i}\log_{2}p_{i}$ (38)
+For ANNs, continuous activations result in higher entropy, enabling richer information flow.
+In contrast, SNNs transmit information only when neurons spike, which limits entropy and reduces the system’s capacity to capture detailed patterns[3].
+Despite its efficiency, MK-SGN’s accuracy is constrained by this sparsity.
+It outperforms other SNN models but still lags behind ANN-based models in high-precision tasks.
+The trade-off between energy efficiency and accuracy remains a challenge.
+Application Scenarios
+While SNNs offer exceptional energy efficiency, their accuracy is often lower than that of GCNs, especially in tasks requiring detailed pattern recognition.
+SNNs are particularly suited for real-time applications where energy efficiency is a priority.
+For example, in wearable health monitoring systems, SNNs can be used for continuous gesture and activity recognition with relatively low energy consumption, making them ideal for long-term, battery-powered use.
+Similarly, in mobile robotics, SNNs can process sensor data efficiently, enabling real-time action recognition and decision-making on low-power devices, such as drones or robotic assistants.
+These applications require real-time, low-energy operation, where the trade-off between accuracy and energy efficiency is critical.
+In contrast, GCNs excel in tasks requiring high accuracy and detailed spatial and temporal relationships, such as skeleton-based action recognition in complex environments.
+GCNs are ideal for scenarios where precise behavior analysis is necessary, such as in surveillance systems for human activity recognition in security or healthcare settings.
+For example, in smart home monitoring systems, GCNs can analyze detailed human motion patterns to detect falls, abnormal behaviors, or to improve user experience through activity-based control.
+Similarly, in sports analytics, GCNs can be used to analyze player movements in a game, providing insights into strategies, player health, and performance.
+These scenarios require high accuracy to capture the fine-grained details of human motion, which GCNs are well-suited for.
+Future Work
+To enhance the performance of MK-SGN, future work should focus on optimizing spike coding, increasing spike rates, or exploring hybrid architectures with ANNs, like GCNs or CNNs, to enhance both accuracy and efficiency.
+SNNs can benefit from advanced spike encoding schemes that allow for richer information transmission while maintaining low energy consumption.
+Hybrid models that combine the strengths of SNNs and GCNs can offer a balance between real-time energy efficiency and high accuracy, especially in complex action recognition tasks.
+Additionally, training methods tailored to SNNs should be developed to improve performance in high-accuracy tasks.
+Approaches such as spike-timing dependent plasticity (STDP) or supervised learning algorithms specifically designed for spiking neurons could lead to better convergence and higher recognition accuracy in the context of skeleton-based action recognition.
+These advancements will help address the current limitations and broaden the application of SNN-based models in real-world scenarios.
 5 Conclusion
-In this work, we propose the Spiking Graph Convolutional Network (MK-SGN), a novel architecture that combines the strengths of SNNs and graph convolutional networks GCNs for skeleton-based action recognition.
-This framework is, to the best of our knowledge, the first to effectively integrate spiking neural computation with graph-based modeling, offering a new perspective on energy-efficient and high-performance recognition methods.
+In this work, we propose the Spiking Graph Convolutional Network (MK-SGN), a novel architecture that combines the strengths of SNNs and GCNs for skeleton-based action recognition.
+To the best of our knowledge, this framework is the first to effectively integrate spiking neural computation with graph-based modeling, offering a new perspective on energy-efficient and high-performance recognition methods.
 The MK-SGN addresses the critical challenge of high energy consumption in traditional GCN-based methods while maintaining competitive performance.
-By leveraging multimodal fusion and knowledge distillation, MK-SGN not only optimizes the use of spiking neural computation but also enhances the ability to extract and integrate meaningful features across modalities.
-These innovations enable MK-SGN to significantly reduce energy consumption on three benchmark skeleton action recognition datasets, achieving over 98% reduction compared to the best ANN-based methods.
-Moreover, MK-SGN consistently surpasses SOTA SNN models in recognition accuracy, achieving an average improvement of 5%.
-This highlights its ability to balance energy efficiency and computational effectiveness, making it particularly well-suited for real-time and resource-constrained applications.
-These results establish MK-SGN as a robust and practical solution for skeleton-based action recognition, paving the way for future exploration of spiking-based multimodal frameworks in other domains.
+Leveraging multimodal fusion and knowledge distillation, MK-SGN optimizes spiking neural computation and enhances the ability to extract and integrate meaningful features across modalities.
+These innovations enable MK-SGN to substantially decrease energy consumption across three benchmark datasets for skeleton action recognition, achieving an impressive reduction of over 98% compared to the most effective ANN-based methods.
+Moreover, MK-SGN consistently outperforms SOTA SNN models in recognition accuracy, achieving an average improvement of 5%.
+This demonstrates its ability to balance energy consumption with recognition performance, making it well-suited for resource-constrained applications.However, the accuracy of MK-SGN is still limited by the sparsity inherent in SNNs.
+While the model achieves significant energy efficiency, its performance in high-precision tasks still lags behind traditional ANN-based methods, primarily due to the sparse information transfer characteristic of spiking networks.
+Future work should focus on addressing these limitations by exploring advanced spike coding techniques to improve information transmission and enhance accuracy.
+Additionally, incorporating hybrid models that combine the strengths of SNNs and ANNs could help balance energy efficiency and accuracy more effectively.
+Optimizing spike rates and developing tailored training methods for SNNs, could also boost performance in high-accuracy scenarios.
+Despite these challenges, MK-SGN offers a promising approach for energy-efficient skeleton-based action recognition, and its potential for real-time applications in resource-constrained environments positions it as a solid foundation for future exploration in spiking-based multimodal frameworks.
+6 Acknowledgements
+The National Natural Science Foundation of China NSFC-61976022 supports this work.